A Bibliography of Publications of Nicholas Constantine Metropolis

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

12 October 2024 Version 1.109

Abstract

This bibliography records publications of Nicholas Constantine Metropolis (1915–1999).

Title word cross-reference

(0, 1) [EM72, MS67, MSS69]. 10 [S.69]. 15 [BMSW54, S.69]. 16 [BMSW54]. 2 [S.69]. 2,000 [MRvN50]. 2600 [Met41a, Met41c]. $29.50 [Tro83]. 3 [Fan60, RBMW59, Sho60]. 6 [Fan60, RBMW59, Sho60]. 88062 [S.69]. ^{3} [KKHM54]. ^{4} [KKHM54]. _{2} [Met41b]. e [MRvN50, Rei50]. j [Fan60, RBMW59, Sho60]. λ3400 [Met41a, Met41c]. N [BL84, EM71a, LM81, MR78a, MR78b]. ν [EM71a]. p [EM75]. π [MRvN50, Rei50]. Q_{p} [EM75]. X_{m} + 1 [EM75].

-Adic [EM75]. -City [BL84]. -cube [MR78a, MR78b]. -dimensional [LM81]. -matrices [EM72].

12th [Jac77]. 12th-15th [Jac77]. 15th [Jac77, Jun64a, Jun64b]. 16-19 [MTTT63a]. 175 [FNR^{+}07]. 1954 [Mey56]. 1962 [MTTT63a]. 1965 [Kal65]. 1972 [IEE72b]. 1976 [Jac77]. 1983 [Gre84]. 1985 [A^{+}85, GM86]. 1995 [ACM95].

20th [Cip00]. 24 [Sha71]. 2nd [IEE72a].

310-Mev [SYM57]. 35 [Met65d]. 3rd [IEE75].

40th [MKR87].

65th [Nac81]. 6600 [GM70b].

810A [MPF66].

A-bomb [Rav99]. A.M.S [Jun64a, Jun64b]. Absorption [Met41c, Met41a]. Academic [Tro83]. Acceptance [Rob99]. ad [Seg86a]. adaptive [CS05, HST01]. Addition [FMRR77b, FMRR77a]. Adic [EM75]. Adventures [Met76a]. Against [TMRL89]. Age [Met92, Met93]. ai [Seg83]. al [LTC^{+}81, LB81, SR82]. Alamos [A^{+}85, GM86, MOR76, MKR87, Ano96, HPR14, Mac91, MN82, Met87b, Met90]. algebra [Gre84, MR74, MR83]. algebraic [RDM62]. Algirdas [Met65d]. Algorithm [BS00, BD01, BL84, CG95, DN04, Hit03, Sas91, WW05, Cal05, CS05, DSC95, DJ20, FNR^{+}06, FNR^{+}07, Gub05, Gus98, HST01, Hol98, MRT73, Ros03, Van17, DSC98]. Algorithms [Cip00, FM68, MM67, Met65a, Met68, RT96, Rob99, BJ24, MT96, Met65c]. Almost [Tau61b]. American [MTTT63a]. Amounts [KMTW43]. AMS [Gre84]. AMS-NSF [Gre84]. Analysis [FMRR77b, FMA54, HMW55, Jac77, Met41c, Met65b, SYM57, Tau63b, dHMAB54, FMRR77a, Met41a, Nac81]. Analyzed [Met72a, Met72b, Met72c, Met73]. Anecdotes [TMRL89]. Anniversary [MKR87]. annual [ACM95, IEE72b]. aos [Seg87]. Application [BM75, BM77, BBL^{+}93, Cal05, ITC74]. Applications [Has70, A^{+}85, Nac81]. Applied [BMN70a, HBLM96, Jun64a, Jun64b, MTTT63a, MOR76]. approach [BMP72, BMP73, GM70a]. Approximation [EM71a, Met41b]. April [A^{+}85, Jac77, MTTT63a]. architecture [IEE72b]. Arithmetic [AM59, BM75, BM77, GM70b, Ham65a, Ham65b, IEE72a, IEE75, Jun64a, Jun64b, MA58, MA63, Met65b, MA65, BM68, FM68, GM70a, MM67, MTTT63a, Met65a, Met65c, Met68, MRT73, MR74, Met76c, Met77, MT77, Met65d]. arithmetical [MU52, MU53]. Art [Jac77]. Astrophysics [Tau63c]. Atlantic [MTTT63a]. atomic [MR50, MR51a]. atomique [A^{+}85]. Automata [Tau63b]. Automatic [TMRL89]. aux [Seg84]. Avisienis [Met65d].

B [Ham65a, Ham65b, Jun64a, Jun64b]. Band [Met39]. Bands [Met41b, Met41a]. Based [FMT49b, FMT49a]. Bases [BMN70a, BMN70b, S.69, Sha71]. Basic [MA63]. Beginning [Met87a, Ano96, Met85]. Behind [Ano96]. Ben [Edw85]. Berichtigung [Met65c]. Berkeley [Met50]. Best [Cip00]. Bets [HPR14]. between [MTTT63b, dHMAB54]. Beyer [S.69, Sha71]. Bibliographic [Yos02]. Big [Met78, Ano78]. bijective [MR88]. Binary [Met65d, Met72a, Met72b, Met73, S.69, Met72c, MR74]. Binary-Compatible [Met65d]. Biographies [Cor87, Wei88]. birthday [Nac81]. Bivins [Fan60, Sho60]. bomb [Rav99]. Book [Edw85, Fan60, Ham65a, Ham65b, Jun64a, Jun64b, Met50, Met76a, S.69, Sho60, Ste81, Tro83]. Boolean [FMRR77a, FMRR77b]. Boulder [Gre84]. Brains [Met50]. Braithwaite [Edw85].

C [Edw85, Ham65a, Ham65b, Jun64a, Jun64b, Met50]. C. [LTC^{+}81]. Cadarache [A^{+}85]. calculated [MRvN50]. Calculation [AM65]. Calculations [Bar65, FM47, FMT50, MRR^{+}53, MRR^{+}52, MBS^{+}58b, MBS^{+}58a]. California [IEE72b]. candidate [CS05]. Capsule [LTC^{+}81, LB81, SR82]. Carlo [Bra83, Edw85, FNR^{+}07, GM86, Ste81, Tro83, A^{+}85, And86, Bar65, BHMU85, DH87, Dys12a, FNR^{+}06, HPR14, Has70, HMG55, HMG56, Hur85, Mar56, MU49b, MBS^{+}58b, MBS^{+}58a, Met85, Met87a, Mey56, Pes73, RC04, RR54, Ros03, TFY95]. carrying [MRT73, MT77]. Cascade [MTB59]. cascades [MBS^{+}58b, MBS^{+}58a]. Case [TMRL89, Van17]. Castle [A^{+}85]. cathedral [Dys12b]. Cave [Edw85]. CDC [GM70b]. Cell [HMG55, HMG56]. Central [RT96]. Century [Lid01, Cip00, Har10, LTC^{+}81, LB81, MHR80, SR82, Ste81, Bra83, Tro83]. certain [GLMU56]. Chains [GM54, Has70, Pes73, Tie94, MRSW78]. Characters [BMSW54]. Chicago [MTTT63a]. Chronicle [Lid01]. City [BL84, Kal65, MTTT63a]. class [MS67, MRSW78, MNR81]. classes [MR91, MRS91, MRS95]. Classical [LM86a]. Classification [BLMS91, BLMS96]. closer [Har10]. CLTs [Rob99]. Codes [HMW55]. coefficients [KKHM54]. Collected [Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c]. Collection [MHR80, KBM^{+}84]. Collector [Gre55]. College [IEE72a]. Colorado [Gre84]. Combinatorial [MR78a]. Combinatorics [Gre84]. Comments [BHMU85]. Commissariat [A^{+}85]. Communications [FNR^{+}07]. Comp [Sha71]. Comparison [RU18]. Compatible [Met65d]. COMPCON [IEE72b]. complete [BLMS96]. completely [EM72]. complicated [May76]. Compounds [KMTW43]. comprehensive [GM70a]. Computation [MR93, SS05]. computational [Cal05]. Computations [BHMU85, Hur85]. Computer [AFI72, AM65, CMR^{+}83, Edw85, FNR^{+}07, HM83, HMG56, IEE72b, IEE72a, IEE75, MA58, MPF66, GM70a, ITC74]. Computers [Edw85, HM83, Rav99, Rot86, Tau63b, KBM^{+}84, Met60]. Computing [ACM95, Bra83, Ham65a, Ham65b, HM83, Jun64a, Jun64b, LTC^{+}81, LB81, MRR^{+}53, MTTT63b, Met72a, Met72b, Met73, MHR80, MW80, MN82, Met92, Nas90, SR82, Ste81, Yos02, MRR^{+}52, MTTT63a, Met72c, MW72, Met93, RM49, Tro83, LTC^{+}81]. Conference [AFI72, Gre84, GM86, IEE72b, Jac77]. congress [Kal65]. Constantine [BBLS00]. constants [Lia10]. contemporanea [Seg76, Seg83]. contemporary [Seg76]. Continued [RDM62]. Continuous [Tau62]. Convergence [BJ24, RT96, Hol98, MT96]. Conversion [MA65]. Coordinate [MTB59]. Corrigendum [Sha71]. coupled [BMP72, BMP73]. Coupon [Gre55]. Crises [Edw85]. Criteria [Rob99]. Critical [KMTW43]. cube [MR78a, MR78b]. cycles [BLMS91, BLMS96]. cyclic [MSS69]. cyclotomic [MR84].

dai [Seg83]. Dallas [IEE75]. Data [Cor87]. David [Edw85]. Decimal [Rei50, MRvN50]. Decomposition [Met65b, Met68]. découvertes [Seg84]. Dedicated [Rot86, MOR76, Nac81]. Defense [BCMR92]. defined [GLMU56]. degeneracy [LM81]. Degree [BMSW54]. Demons [Dys12c]. Density [Sas91, CS05]. descobertas [Seg87]. Design [Tau63b]. determinants [MS67]. Determination [Rei50]. Deuterium [FMT50]. Deuterium-Tritium [FMT50]. Development [HBLM96, Mac91]. Dictionary [Cor87]. Differential [BM75, BM77]. diffusion [RUvN47]. digest [IEE72b]. Digit [GM70b, MA58, Met65d]. Digital [HMG56, Dys12b]. Digits [BMN70b, Gre55, S.69, MRvN50, Sha71]. Dimensional [BMv^{+}59b, BMv^{+}59a, LM81]. Dioxide [Met41c, Met41a]. Dirac [MR51b]. Directions [MKR87]. discoveries [Seg76, Seg80, Seg84, Seg87, Seg07]. distributed [MU49a]. Distribution [Bar65]. Distributions [Tie94]. DNA [BBL^{+}93]. Do [DSC98, DSC95]. Donald [Edw85]. double [Lia10]. drop [FM47]. during [Gre84]. Dynamics [LM86b, LM86a, May76].

Early [BHMU85, Hur85, MN82]. ed [Bra83]. Edited [Tro83]. Editors [Cip00]. Edmund [Met50]. Edward [Har10]. Effect [Sas91]. Electron [Bar65]. Electronic [Met41b, Met41a]. elementary [MS70]. Elements [FMT49b, FMT49a]. énergie [A^{+}85]. energy [LM80, LM81, MBS^{+}58b, MBS^{+}58a]. ENIAC [FMT50, HPR14, MRvN50, Rei50]. Enrico [Met54, Met55]. enumeration [MS80]. Equation [BM75, BM77, MRR^{+}52, MRR^{+}53, BMP72, BMP73, MR51b]. Equations [FMT49a, FMT49b, Met56, RR54]. Equivalent [S.69]. Era [HM83, MR93]. Ergodic [Tau61b]. Erratum [FNR^{+}07]. Error [AM65]. Errors [Met65b, MW80, MW72]. Essays [HR00, MOR76, MHR80, Wat04b, KBM^{+}84, Nac81]. Estimation [AM65]. evaluation [MM67, Met68]. Evans [LTC^{+}81]. every [Van17]. Excellence [Lid01]. Expansions [BMN70a, RDM62]. expectations [Met85]. Experience [Met87b, Met90]. experiences [Ano96]. Experimental [Ham65b, Jun64a, Jun64b, MTTT63a, Ham65a]. Experiments [SYM57]. Exploring [Tie94]. Extension [WW05].

fa [Seg86b]. faces [MR78a, MR78b]. Family [Edw85]. Fast [MRR^{+}53, MRR^{+}52]. fences [Ano96]. Fermi [FMT49a, FMT49b, MR51b, Met54, Met55]. Field [EM75]. fifteenth [MTTT63a]. fifty [DJ20]. finite [MSS73]. first [MRvN50, MOR76, AFI72]. fisica [Seg76, Seg83]. fisicos [Seg87]. fisikaim [Seg86a]. fission [FM47]. Floating [AM59]. Florida [Mey56]. Forester [SR82]. four [BBL^{+}93]. fraction [RDM62]. France [A^{+}85]. Francisco [IEE72b]. French [Seg84]. Frontiers [GM86, MSWA86]. function [MU52, MU53]. Functions [Bar65, Tau61b, MNR81, MR88, MR91, MRS95, BL01]. Further [RR54]. Future [HR00].

G [LTC^{+}81]. Games [Tau63c]. Gauss [EM71b]. generalization [EM71b]. Generalized [BMN70a, FMT49b, FMT49a]. generators [Cal05]. Genesis [Ros03]. Geometric [BD01, Hol98, RT96]. Geometry [Tau62]. Gian [Bra83, Edw85, Ste81, Tro83]. Gian-Carlo [Bra83, Edw85, Ste81, Tro83]. Giant [Met50]. Gibbs [Van17]. Gilbert [Edw85]. graphical [MS80]. Great [Met78, Ano78, Met85]. Greek [Wat04a]. Group [HBLM96, Tau61b]. Groups [BMSW54, LM86a]. Guide [Yos02]. guided [Gus98].

H [Edw85, Ham65a, Ham65b, Jun64a, Jun64b]. Handbook [BL01]. harmonic [LM80, LM81]. Hastings [BD01, BJ24, CG95, CS05, DJ20, Hit03, Hol98, Lia10, MT96, RT96, Rob99, Van17]. held [A^{+}85, Gre84, Jac77, MTTT63a, Mey56]. Hidden [LM80]. High [Ham65a, Ham65b, Jun64a, Jun64b, MTTT63b, Met60, MTTT63a, MBS^{+}58a]. High-energy [MBS^{+}58a]. High-Speed [MTTT63b]. Historical [Cor87]. History [Bra83, Hit03, LTC^{+}81, LB81, MHR80, MW80, Nas90, SR82, Ste81, MW72, MA87, Tro83]. hit [RU18]. hit-and-run [RU18]. honor [HR00]. Hotel [IEE72b]. Howlett [Bra83, Ste81, Tro83]. Hydrodynamical [BMv^{+}59b, BMv^{+}59a]. Hydrodynamics [Tau63c]. Hydrogen [FMA54, dHMAB54].

ICR [Met64b, Met64a, Met65e]. identity [MR84, MR88]. IEEE [IEE72b, IEE72a]. IEEE-TCCA [IEE72a]. IFIP [Kal65]. Ignition [FMT50]. II [MPF66, LMO^{+}56, MM67, Met41a, MBS^{+}58a, Tau61b]. III [FM68, Tau63a]. Illinois [MTTT63a]. imprecise [Met79, Met80b]. Incremental [Edw85]. independent [CS05]. infinite [MU49a]. Influence [Mac91]. influential [Har10]. Information [Edw85, KBM^{+}84, Kal65]. Inherent [Met65b]. Innovative [IEE72b]. Integers [BMN70b, S.69, GLMU56, Sha71]. Interactions [MTTT63b]. International [IEE72b, ITC74]. Interpretation [BD01]. interval [MSS73]. interview [MA87]. intractable [Lia10]. Intranuclear [MTB59, MBS^{+}58b, MBS^{+}58a]. introduction [Cal05]. introductory [Mar56]. Inventos [LTC^{+}81]. inversion [FM68]. Iodide [Met39]. Irrational [BMN70a]. isotropic [LM81]. Italian [Seg76]. iterated [EM71b]. IV [Tau62].

J [Bra83, S.69, Sha71, Ste81, Tro83]. J. [LTC^{+}81]. Jack [IEE72b]. Japan [AFI72, AFI72]. Jersey [MTTT63a]. John [Fan60, Ham65a, Ham65b, Jun64a, Jun64b, Sho60, CMR^{+}83, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c]. Joint [Gre84, A^{+}85]. Jr. [Edw85, Fan60, Sho60]. Judging [Har10]. June [ACM95, Gre84].

Karl [Edw85]. karne [Seg86a]. Kerr [Edw85]. Know [DSC98, DSC95]. kvarkim [Seg86a].

L [SR82]. Laboratories [Mac91]. Laboratory [A^{+}85, GM86]. Laboratory-Commissariat [A^{+}85]. Labs [BCMR92]. Languages [SR82]. lattice [MR78b]. lattices [TFY95]. Laurent [Nac81]. Leads [HM83]. Leap [BCMR92]. Letter [BCMR92]. letters [BBL^{+}93]. leurs [Seg84]. levels [LM80, LM81]. li [Seg86b]. Lie [LM86a]. Lifecycle [Edw85]. Limit [RT96, EM71b, MSS73]. line [MU49a]. linear [MSS67]. liquid [FM47]. liquid-drop [FM47]. Livermore [Mac91]. Logic [Tau61a]. logical [Met60]. look [Har10]. Low [MBS^{+}58b]. Low-energy [MBS^{+}58b].

M [Edw85, Met76a, MOR76]. Machines [Met50, MRR^{+}53, MRR^{+}52]. Maker [Rav99]. Management [Edw85]. MANIAC [And86, Ano78, DJK^{+}53, HMW55, JM51, JM54, LMO^{+}56, MPF66, Met80a, Met78]. Manuel [Fan60, Sho60]. Many [Tel55]. map [BLMS91, BLMS96]. Maps [LM86b]. March [Mey56]. Markov [Has70, Pes73, Tie94]. Marshall [Gub05]. Maryland [IEE72a, IEE72a]. mass [MU49a]. masses [MR50, MR51a]. Math [Sha71]. Mathematical [BL01, HBLM96, MTTT63b, MTTT63a, Nac81, May76]. Mathematician [Met76a]. Mathematics [HBLM96, Jun64a, Jun64b, MTTT63a, MOR76, Ham65b, Jun64b, Ham65a, Jun64a]. matrices [EM72, MS67, MSS69]. Matrix [Met65b, Met68, FM68]. May [ACM95, IEE72a, Kal65]. Maymon [Edw85]. MCNP [DH87]. means [EM71b]. Measurements [Lid01]. Mechanics [BL84, Tau61a, Cal05, Ros03]. Meeting [A^{+}85]. Meetings [BHMU85, Hur85]. Memoir [Met92, Met93]. Memoirs [Met76b]. Meteorology [Tau63c]. Method [DH87, Met87a, FNR^{+}06, FNR^{+}07, Mar56, MU49b, TFY95, Edw85]. Methodist [IEE75]. Methods [Has70, HMG55, HMG56, Met76c, Met77, Mey56, A^{+}85, RUvN47, RC04]. Metropolis [Bra83, Edw85, Fan60, FNR^{+}07, Ham65a, Ham65b, Jun64a, Jun64b, LTC^{+}81, LB81, S.69, Sha71, Sho60, SR82, Ste81, Tro83, And86, Ano00, BBLS00, BS00, BD01, BL84, BJ24, Cal05, CG95, CS05, DSC95, DSC98, DN04, DH87, FNR^{+}06, Gub05, Gus98, HST01, HR00, Hei99, Hit03, Hol98, Lia10, MT96, MA87, Rav99, RT96, Rob99, Rot86, RU18, Sas91, TFY95, Van17, WW05]. Mev [SYM57, dHMAB54]. Mexico [GM86, Wat04b]. Mi [Seg86a]. Mi-karne [Seg86a]. Microelectronics [SR82]. middle [Met56]. ming [Seg86b]. Minimum [FM52]. Mixtures [FMT50]. Modal [BMP72, BMP73]. model [FM47]. models [Lia10, May76]. Modern [RM49, Seg80, Seg87, Seg07, Seg84]. modernes [Seg84]. modernim [Seg86a]. modernos [Seg87]. Molecules [Met41b, Met41a]. Monte [FNR^{+}07, GM86, A^{+}85, And86, Bar65, BHMU85, DH87, Dys12a, FNR^{+}06, HPR14, Has70, HMG55, HMG56, Hur85, Mar56, MU49b, MBS^{+}58b, MBS^{+}58a, Met85, Met87a, Mey56, Pes73, RC04, RR54, Ros03, TFY95]. Monte-Carlo [A^{+}85]. most [Har10]. motions [MU49a]. Multidimensional [RT96]. Multiplication [FMRR77b, HMG56, FMRR77a]. My [Edw85].

N [Bra83, Edw85, Fan60, Ham65a, Ham65b, Jun64a, Jun64b, S.69, Sha71, Sho60, Ste81, Tro83]. N. [LTC^{+}81, LB81, SR82]. Name [Cip00]. National [A^{+}85, GM86, Mac91]. Natural [MOR76]. NBS [Lid01]. NBS/NIST [Lid01]. necklaces [MR83]. Neergaard [S.69, Sha71]. Negative [FMA54]. nella [Seg76, Seg83]. Neumann [Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c]. neutron [RUvN47]. neutronics [A^{+}85]. Nevada [ACM95]. News [Ano00]. Nicholas [Ano00, BBLS00, MA87, Rav99, Rot86]. Nick [HR00, Hei99, Wat04a]. NIST [Lid01]. non [MSS67]. non-linear [MSS67]. Nonlinear [BMP72, BMP73]. normal [EM72]. normalized [Met68]. normalizing [Lia10]. Note [Hur85, Rob99, TFY95, Mar56]. November [IEE75]. NSF [Gre84]. Nuclear [HPR14]. Number [LM81, Cal05, LM80, Met64b, Met65e]. Number-theoretic [LM81, LM80]. numbers [FMRR75, FMRR76, Met79, Met80b, RDM62]. Numerical [BMv^{+}59b, DN04, FM52, FMT50, Jac77, Tau63b, BMv^{+}59a, SS05]. Numerology [GM54].

Obituary [BBLS00, Hei99, Rav99]. Obra [LTC^{+}81]. occasion [Nac81]. October [AFI72]. offprints [Met54]. Oh [Wei88]. one [Har10]. onslaught [KBM^{+}84, Edw85]. Operations [MA63]. Operators [Tau63a, Tau61b]. Optimum [Pes73]. Oral [MA87]. ordered [MRSW78]. organization [Met60]. Originating [HBLM96]. origins [Dys12b]. oscillator [LM80, LM81]. oscillators [BMP72, BMP73]. Other [Edw85, Tau62]. ove [Seg86a]. ove-ad [Seg86a]. ove-tagliyotehe [Seg86a]. Own [BCMR92].

papers [IEE72b, MTTT63b]. parabolic [BLMS91, BLMS96]. Park [IEE72a]. Partial [BM75, BM77]. partially [MRSW78]. Partitions [MRSW78, MS70, MS80]. Pasta [CMR^{+}83]. Path [CMR^{+}83]. People [Tel55]. perfect [CS05]. Periodic [Tau61b]. Permanents [MSS69]. Personaggi [Seg76, Seg83]. Personal [Edw85, Met92, Met93].

Personalities [Seg76]. perturbed [LM80, LM81]. Phase [FMA54, Met56, SYM57, dHMAB54]. Phase-Shift [SYM57]. photonics [A^{+}85]. Physical [HBLM96]. physiciens [Seg84]. physicists [Seg80, Seg84, Seg87, Seg07]. Physics [FNR^{+}07, MKR87, A^{+}85, Cal05, Seg76]. Pion [dHMAB54, MBS^{+}58a]. Pion-Hydrogen [dHMAB54]. Pioneers [LTC^{+}81, Wei88]. Pions [FMA54]. Places [Rei50]. Plasma [Bar65]. Point [AM59]. points [MU49a]. Polyatomic [Met41b, Met41a]. Polynomial [Met68, MM67]. Polypeptide [GM54]. Populations [HMG55]. Portuguese [Seg87]. Posterior [Tie94]. pp [Tro83]. Precise [SS05]. Press [Tro83]. probability [MT77]. Problem [BMv^{+}59b, BL84, FM52, HMW55, BMv^{+}59a, MS70]. Problems [HBLM96, Cal05]. Proceedings [AFI72, Jac77, Jun64b, MTTT63a, A^{+}85, Kal65, MOR76, ACM95, Gre84, GM86, Jun64a]. processes [MBS^{+}58a]. Processing [Cor87, Kal65]. product [FMRR75, FMRR76]. Programming [SR82, TMRL89]. Prolate [Met41b]. Properties [BBL^{+}93]. property [MU52, MU53]. Proton [Bar65, SYM57]. Proton-Electron [Bar65]. Proton-Proton [SYM57]. Provence [A^{+}85]. Pseudo [CS05]. Pseudo-perfect [CS05]. Publications [Lid01].

Quantum [GM86, Tau61a]. quark [Seg83]. quarks [Seg80, Seg84, Seg87, Seg07]. Quarterly [Met64b, Met64a, Met65e]. Quevedo [LTC^{+}81].

R [CMR^{+}83, Fan60, S.69, Sha71, Sho60]. R. [SR82]. R65 [Met65d]. R65-35 [Met65d]. Radial [Bar65]. Radix [MA65]. raggi [Seg83]. raios [Seg87]. Random [Cal05, Gre55, Met96, RU18, TFY95]. randomly [MU49a]. randomness [MU52, MU53]. Ranotgen [Seg86a]. Rate [Rob99]. Rates [MT96, BJ24]. rayons [Seg84]. rays [Seg80, Seg84, Seg87, Seg07]. real [FMRR75, FMRR76]. Recurrence [Met65c, Met65a]. relations [Met65a, Met65c]. Reminiscences [Met96]. ren [Seg86b]. Report [Met64b, Met64a, Met65e]. Research [Gre84, MTTT63b]. Resources [Yos02]. restricted [MS70]. Results [DN04, RR54]. Review [Fan60, Ham65a, Ham65b, Jun64a, Jun64b, Met50, Met65d, Met76a, S.69, Sho60, Ste81, Tro83, Bra83]. Reviews [Edw85, LTC^{+}81, LB81, SR82]. Revolution [SR82]. Rings [Tau63a]. root [EM71a]. Roots [BMN70a, BMN70b, EM75, S.69, Sha71]. Rosenbluth [Gub05]. Ross [Edw85]. Rota [Tro83, Bra83, Edw85, Ste81]. Rotenberg [Fan60, Sho60]. run [RU18].

S [Met76a, MOR76]. Salesman [BL84]. sampler [Lia10, RU18, Van17]. Sampling [Has70, Pes73]. San [IEE72b]. Santesmases [LTC^{+}81]. saving [FNR^{+}06, FNR^{+}07]. Scattering [FMA54, SYM57]. Schneider [Edw85]. Schwartz [Nac81]. Science [CMR^{+}83, Edw85, KBM^{+}84, Rot86, ITC74]. Sciences [MOR76]. Scientific [BHMU85, Hur85, Nas90, Yos02]. scientists [Har10]. scoperte [Seg76, Seg83]. Second [KKHM54]. SEL [MPF66]. SEL-810A [MPF66]. Selected [Lid01, MTTT63b]. Selma [MPF66]. sense [OMRS84]. September [GM86, IEE72b]. sequences [BBL^{+}93, GLMU56]. Serial [BMN70a]. Sets [Tau61a, MSS73, MRSW78]. seventh [ACM95]. Sharp [Edw85]. Shift [FMA54, SYM57, dHMAB54]. shifts [Met56]. sieves [GLMU56]. Signed [Met65d]. Signed-Digit [Met65d]. Significance [BM75, BM77, MRT73, MR74, MT77, Met76c, Met77]. Significant [GM70b, MA58]. Silver [Met39]. Simple [May76]. Simulation [HM83, MPF66, BM68, Hol98]. Simulations [HPR14]. Skiing [Wat04b]. slice [RU18]. Society [IEE72b, MTTT63a]. Software [Edw85]. Solution [BMv^{+}59b, FM52, BLMS96, BMv^{+}59a, MS70]. Solutions [MR51b]. Some [BMN70a, BMN70b, Met85, Sha71]. Southern [IEE75]. spatial [Lia10]. special [Van17]. Speed [Ham65a, Ham65b, Jun64a, Jun64b, MTTT63b, Met60, MTTT63a]. Square [BMN70a, BMN70b, S.69, Sha71]. squares [Met56]. Stable [MSS67]. Standards [Lid01]. State [FMT49b, Jac77, MRR^{+}53, FMT49a, MRR^{+}52, RR54]. States [Sas91, MSS67]. Statistical [BMN70b, BL84, MRvN50, RUvN47, Sha71, A^{+}85, RC04, Ros03]. stories [Ano96]. strings [MR74]. Structure [Met41b, LM80, Met41a, MR78a]. structures [BM68]. Studies [HMG56, MTB59, MBS^{+}58b, MBS^{+}58a]. Study [BMN70b, BMv^{+}59b, HMG55, BMv^{+}59a, Sha71]. suas [Seg87]. Sulphur [Met41c, Met41a]. Summation [Met79, Met80b]. Summer [Gre84]. Sun [Wat04b]. Supercomputing [Mac91, MSWA86]. Surveys [MOR76]. Symbolic [LM86b]. symbols [RBMW59, Fan60, Sho60]. Symmetric [BMSW54, MR88, MNR81]. symmetrizability [EM72]. Symmetry [MR91, MRS95, LM80, MRS91]. Symposia [MTTT63b]. Symposium [ACM95, IEE72a, IEE75, Jun64a, Jun64b, Met60, MTTT63a, MOR76, Mey56]. System [Met39, Met41c, MA63, MA65, Met41a]. systems [MU49a].

T. [SR82]. Table [MR50, MR51a]. tagliyotehe [Seg86a]. tall [Ano96]. Tar [IEE72b]. Taub [Ham65a, Ham65b, Jun64a, Jun64b]. TCCA [IEE72a]. Technology [Lid01, ITC74]. Teller [Har10]. Test [BMN70a, Gre55]. Texas [IEE75]. th [EM71a]. Their [Has70, ITC74, Seg80, Seg84, Seg87, Seg07]. Theorems [RT96]. theoretic [LM80, LM81]. Theory [ACM95, FMT49b, HBLM96, MRS91, Tau61a, Tau61b, Tau63b, FMT49a, Tau63c]. Think [Met50]. Thomas [MR51b, FMT49a, FMT49b]. those [BBL^{+}93]. three [MR91]. time [FNR^{+}06, FNR^{+}07]. Todd [Ham65a, Ham65b, Jun64a, Jun64b]. Tokyo [AFI72]. Tompkins [Jun64a, Ham65a, Ham65b, Jun64b]. Top [Cip00]. Topics [Tau62]. Torres [LTC^{+}81]. Toy [Met78, Ano78]. tracts [ITC74]. transformation [MSS67]. Transformations [MTB59, MSS73]. Trapezoidal [LM86b]. Travelling [BL84]. treatment [MRvN50]. Trilogy [MW80, MW72]. Tritium [FMT50]. Tumor [HMG55]. Turing [Dys12b]. Twentieth [Bra83, LTC^{+}81, LB81, SR82, Ste81, Tro83, Har10, MHR80]. twenty [ACM95]. twenty-seventh [ACM95]. Two [BMv^{+}59b, BBL^{+}93, BMv^{+}59a, BM68]. Two-Dimensional [BMv^{+}59b, BMv^{+}59a].

Ulam [MOR76, Dys12c, Met76a]. Ultraviolet [Met39]. un-normalized [Met68]. Uncommon [OMRS84]. Understanding [CG95]. unit [MSS73]. universe [Dys12b]. University [Gre84, IEE75, Jac77, Mey56]. Unnormalized [AM59, MA63, Met65b, MA65, FM68, MM67, Met65a, Met65c]. Unrestricted [MM67]. Unusual [CMR^{+}83]. Uranium [KMTW43]. USA [AFI72]. USA-Japan [AFI72]. Using [FMT50, Has70, Met65b, Pes73].

V [Tau63b]. values [MRvN50]. vanishing [MS67]. variables [MR91]. variants [CS05]. Various [BMN70a, BMN70b, S.69, Cal05, Sha71]. vectors [MR83]. Vegas [ACM95]. very [May76, Met60]. VI [Tau63c]. Vibrational [Met41c, Met41a]. virial [KKHM54]. Volume [Rot86, MKR87, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c].

W [Edw85, S.69, Sha71]. walk [Gus98, RU18]. wave [Met56]. Weapons [HM83]. Wexelblat [SR82]. William [Edw85]. Witt [MR83]. Wooten [Fan60, Sho60]. words [BBL^{+}93]. Work [DH87, Tel55]. Works [Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c]. wreath [FMRR75, FMRR76]. Wu [Seg86b].

X [Seg80, Seg83, Seg84, Seg87, Seg07]. X-rays [Seg80, Seg84, Seg87, Seg07]. xian [Seg86b]. xix [Tro83]. XY [Met41b].

York [Jac77, Kal65, Tro83].

References

Alcouffe:1985:MCM

[A^{+}85] Raymond E. Alcouffe et al., editors. Monte-Carlo methods and applications in neutronics, photonics, and statistical physics: proceedings of the joint Los Alamos National Laboratory-Commissariat à l’énergie atomique Meeting held at Cadarache Castle, Provence, France, April 22–26, 1985, volume 240 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-387-16070-1 (paperback). LCCN QC20.7.M65 M65 1985. DM66.00 (West Germany).

ACM:1995:PTS

[ACM95] ACM, editor. Proceedings of the twenty-seventh annual ACM Symposium on Theory of Computing: Las Vegas, Nevada, May 29– June 1, 1995. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995. ACM order no. 508950.

AFIPS:1972:FUJ

[AFI72] AFIPS, editor. First USA-Japan Computer Conference, Proceedings: October 3–5, 1972, Tokyo, Japan. AFIPS Press, Montvale, NJ, USA, 1972. LCCN QA76 .U2 1972.

Ashenhurst:1959:UFP

[AM59] Robert L. Ashenhurst and Nicholas Metropolis. Unnormalized floating point arithmetic. Journal of the ACM, 6(3):415–428, July 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Ashenhurst:1965:EEC

[AM65] R. L. Ashenhurst and N. Metropolis. Error estimation in computer calculation. American Mathematical Monthly, 72(2 (Part 2)):47– 58, 1965. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). Reprinted in Department of Economics, Graduate School of Business, University of Chicago, Center for Mathematical Studies in Business and Economics, number 45.

Anderson:1986:MMC

[And86] Herbert L. Anderson. Metropolis, Monte Carlo, and the MANIAC. Los Alamos Science, 14:96–108, Fall 1986. CODEN LASCDI. ISSN

0273-7116. URL http://library.lanl.gov/cgi-bin/getfile?
00326886.pdf;   http://library.lanl.gov/cgi-bin/getfile?
14-05.pdf;   http://library.lanl.gov/cgi-bin/getfile?number14.
htm;   http://www.osti.gov/energycitations/product.biblio.
jsp?osti_id=5697932&query_id=0. Report LA-UR-85-1202;CONF-

8504110-3.

Anonymous:1978:MGB

[Ano78] Anonymous. The MANIAC: a great big toy. Datamation, 24(8): 80–??, ???? 1978. CODEN DTMNAT. ISSN 0011-6963. Interview with N. C. Metropolis, including brief mention of the roles of Richard Clippinger and John von Neumann in the development of the stored program concept.

Anonymous:1996:BTF

[Ano96] Anonymous. Behind tall fences: stories and experiences about Los Alamos at its beginning. Los Alamos Historical Society, Los Alamos, NM, USA, 1996. ISBN 0-941232-91-3. xiv + 210 pp. LCCN QC773.3.U5 B43 1996.

Anonymous:2000:NNM

[Ano00] Anonymous. News: Nicholas Metropolis, 1915–1999. The Bulletin of the Santa Fe Institute, 15(1):36, Spring 2000. CODEN ???? ISSN 1058-7608.

Barker:1965:MCC

[Bar65] A. A. Barker. Monte Carlo calculations of the radial distribution functions for a proton-electron plasma. Australian Journal of Physics, 18(??):119–133, ???? 1965. CODEN AUJPAS. ISSN 0004-9506 (print), 1446-5582 (electronic).

Bell:1993:PWF

[BBL^{+}93] G. I. Bell, R. L. Bivins, J. D. Louck, N. Metropolis, and M. L. Stein. Properties of words on four letters from those on two letters with an application to DNA sequences. Advances in Applied Mathematics, 14(3):348–367, 1993. ISSN 0196-8858 (print), 1090-2074 (electronic).

Balazs:2000:ONC

[BBLS00] Nandor L. Balazs, John C. Browne, James D. Louck, and Daniel S. Strottman. Obituary: Nicholas Constantine Metropolis. Physics Today, 53(10):100, October 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.aip.org/

pt/vol-53/iss-10/p100.html.

Bethe:1992:LLL

[BCMR92] Hans A. Bethe, George Cowan, Nicholas Metropolis, and Louis Rosen. Letter: Labs leap to own defense. Bulletin of the Atomic Scientists, 48(9):45–46, November 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Billera:2001:GIM

[BD01] Louis J. Billera and Persi Diaconis. A geometric interpretation of the Metropolis–Hastings algorithm. Statistical Science, 16(4): 335–339, November 2001. CODEN STSCEP. ISSN 0883-4237 (print), 2168-8745 (electronic). URL http://projecteuclid.

org/euclid.ss/1015346318.

Burks:1985:CEM

[BHMU85] Arthur W. Burks, Alston S. Householder, N. Metropolis, and S. M. Ulam. Comments on early Monte Carlo computations and scientific meetings. Annals of the History of Computing, 7 (2):147–148, April/June 1985. CODEN AHCOE5. ISSN 0164-

1239. URL http://dlib.computer.org/an/books/an1985/pdf/
a2141.pdf;   http://links.jstor.org/sici?sici=0002-9890%
28196502%2972%3A2%3C47%3AEEICC%3E2.0.CO%3B2-T;   http:
//www.computer.org/annals/an1985/a2141abs.htm.

Brown:2024:CRM

[BJ24] Austin Brown and Galin L. Jones. Convergence rates of Metropolis–Hastings algorithms. WIREs Computational Statistics, 16(5):e70002:1–e70002:??, September–October 2024. CO-DEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

URL https://wires.onlinelibrary.wiley.com/doi/epdf/10.
1002/wics.70002.

Bonomi:1984:CTS

[BL84] Ernesto Bonomi and Jean-Luc L. Lutton. The N-city travelling salesman problem: Statistical mechanics and the Metropolis algorithm. SIAM Review, 26(4):551–568, October 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boisvert:2001:HMF

[BL01] Ronald F. Boisvert and Daniel W. Lozier. Handbook of Mathematical Functions. In Lide [Lid01], pages 135–139.

URL https://nvlpubs.nist.gov/nistpubs/sp958-lide/135-
139.pdf;   https://nvlpubs.nist.gov/nistpubs/sp958-lide/

html/135-139.html. NIST Special Publication.

Bivins:1991:CAC

[BLMS91] R. L. Bivins, J. D. Louck, N. Metropolis, and M. L. Stein. Classification of all cycles of the parabolic map. Physica D, Nonlinear phenomena, 51(1–3):3–27, August 1991. CODEN PDNPDT. ISSN 0167-2789 (print), 1872-8022 (electronic). Nonlinear science: the next decade (Los Alamos, NM, 1990).

Bivins:1996:CAC

[BLMS96] R. L. Bivins, J. D. Louck, N. Metropolis, and M. L. Stein. Classification of all cycles of the parabolic map: the complete solution. Technical report, Los Alamos National Laboratory, Los Alamos, NM, USA, 1996.

Blandford:1968:STA

[BM68] R. C. Blandford and N. Metropolis. The simulation of two arithmetic structures. Technical Report LA-3979, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1968.

Bivins:1975:SAA

[BM75] R. L. Bivins and N. Metropolis. Significance arithmetic: Application to a partial differential equation. In IEEE SCA ’75 [IEE75], pages 64–66. LCCN QA76.6.S919 1975. URL http://www.acsel-

lab.com/arithmetic/arith3/papers/ARITH3_Bivins.pdf. Also

available as Los Alamos Technical Report LA-UR-75-1763 CONF-751103-1.

Bivins:1977:SAA

[BM77] Robert L. Bivins and Nicholas C. Metropolis. Significance arithmetic: Application to a partial differential equation. IEEE Transactions on Computers, C-26(7):639–642, July 1977. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.org/iel5/12/35159/01674896.
pdf?tp=&isnumber=35159&arnumber=1674896&punumber=12;
http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
1674896.

Beyer:1970:GST

[BMN70a] W. A. Beyer, N. Metropolis, and J. R. Neergaard. The generalized serial test applied to expansions of some irrational square roots in various bases. Mathematics of Computation, 24(111):745–747, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842

(electronic). URL http://links.jstor.org/sici?sici=0025-
5718%28197007%2924%3A111%3C745%3ATGSTAT%3E2.0.CO%3B2-2.

Beyer:1970:SSD

[BMN70b] W. A. Beyer, N. Metropolis, and J. R. Neergaard. Statistical study of digits of some square roots of integers in various bases. Mathematics of Computation, 24(110):455–473, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://links.jstor.org/sici?sici=0025-5718%
28197004%2924%3A110%3C455%3ASSODOS%3E2.0.CO%3B2-I. See

corrigendum [Sha71].

Bivins:1972:NCO

[BMP72] R. L. Bivins, N. Metropolis, and John R. Pasta. Nonlinear coupled oscillators: Modal equation approach. Technical Report LA-4934, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, January

1, 1972. 15 pp. URL http://www.osti.gov/energycitations/
product.biblio.jsp?osti_id=4651404&query_id=0.

Bivins:1973:NCO

[BMP73] R. L. Bivins, N. Metropolis, and John R. Pasta. Nonlinear coupled oscillators: Modal equation approach. Journal of computational physics, 12(1):65–87, May 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:/

/www.osti.gov/energycitations/product.biblio.jsp?osti_
id=4457833&query_id=0.

Bivins:1954:CSG

[BMSW54] Robert L. Bivins, N. Metropolis, Paul R. Stein, and Mark B. Wells. Characters of the symmetric groups of degree 15 and 16. Mathematical Tables and Other Aids to Computation, 8 (48):212–216, October 1954. CODEN MTTCAS. ISSN 0891-

6837. URL http://links.jstor.org/sici?sici=0891-6837%
28195410%298%3A48%3C212%3ACOTSGO%3E2.0.CO%3B2-B.

Blair:1959:SNSa

[BMv^{+}59a] A. Blair, N. Metropolis, J. von Neumann, A. H. Taub, and M. Tsingou. A study of a numerical solution to a two-dimensional hydrodynamical problem. Technical Report UIUCDCS-R-1959-302, Department of Computer Science, University of Illinois at Urbana-Champaign, Urbana, Illinois, 1959.

Blair:1959:SNSb

[BMv^{+}59b] A. Blair, N. Metropolis, J. von Neumann, A. H. Taub, and M. Tsingou. A study of a numerical solution to a two-dimensional hydrodynamical problem. Mathematical Tables and Other Aids to Computation, 13(67):145–184, July 1959. CODEN MTTCAS. ISSN 0891-

6837. URL http://links.jstor.org/sici?sici=0891-6837%
28195907%2913%3A67%3C145%3AASOANS%3E2.0.CO%3B2-W. Con-

densation of Los Alamos Scientific Laboratory Report LA-2165. Reprinted in [Tau63b, Paper 17].

Brainerd:1983:HCT

[Bra83] John G. Brainerd. A History of Computing in the Twentieth Century ed. by N. Metropolis; J. Howlett; Gian-Carlo Rota (review). Technology and Culture, 24(1):147–148, January 1983. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/891247/pdf.

Beichl:2000:MA

[BS00] Isabel Beichl and Francis Sullivan. The Metropolis algorithm. Computing in Science and Engineering, 2(1):65–69, January/ February 2000. CODEN CSENFA. ISSN 1521-9615 (print),

1558-366X (electronic). URL http://dlib.computer.org/cs/
books/cs2000/pdf/c1065.pdf;   http://www.computer.org/
cse/cs1999/c1065abs.htm.

Calvayrac:2005:RNG

[Cal05] Florent Calvayrac. Random number generators and the Metropolis algorithm: application to various problems in physics and mechanics as an introduction to computational physics. European Journal of Physics, 26(5):S31, 2005. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic). URL http://stacks.iop.org/

0143-0807/26/i=5/a=S04.

Chib:1995:UMH

[CG95] Siddhartha Chib and Edward Greenberg. Understanding the Metropolis–Hastings algorithm. The American Statistician, 49 (4):327–335, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.

org/stable/2684568.

Cipra:2000:BCE

[Cip00] Barry A. Cipra. The best of the 20th Century: Editors name top 10 algorithms. SIAM News, 33(4):1–2, May 2000. ISSN 0036-1437.

URL https://archive.siam.org/pdf/news/637.pdf.

Curtis:1983:JRP

[CMR^{+}83] Kent K. Curtis, N. C. Metropolis, William G. Rosen, Yoshio Shimamoto, and James N. Snyder. John R. Pasta, 1918–1981 — an unusual path toward computer science. Annals of the History of Computing, 5(3):224–238, July/September 1983. CODEN AH-

COE5. ISSN 0164-1239. URL http://dlib.computer.org/an/
books/an1983/pdf/a3224.pdf;   http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4640708;   http://www.computer.
org/annals/an1983/a3224abs.htm. Foreword by Eric A. Weiss.

Cortada:1987:HDD

[Cor87] James W. Cortada. Historical Dictionary of Data Processing: Biographies. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1987. ISBN 0-313-25651-9 (lib. bdg.). xiii + 321 pp. LCCN QA76.15 .C66 1987.

Corcoran:2005:PPA

[CS05] J. N. Corcoran and U. Schneider. Pseudo-perfect and adaptive variants of the Metropolis–Hastings algorithm with an independent candidate density. Journal of Statistical Computation and Simulation, 75(6):459–475, 2005. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Doolen:1987:MCW

[DH87]  Gary D. Doolen and John Hendricks.  Monte Carlo at

work: MCNP and the Metropolis method. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984):142–143, 1987. CODEN LASCDI. ISSN 0273-7116. URL http://

library.lanl.gov/cgi-bin/getfile?00326867.pdf;   http:
//library.lanl.gov/cgi-bin/getfile?15-12.pdf.

deHoffmann:1954:PHP

[dHMAB54] F. de Hoffmann, N. Metropolis, E. F. Alei, and H. A. Bethe. Pionhydrogen phase shift analysis between 120 and 217 Mev. Physical Review, 95(6):1586–1605, September 15, 1954. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/pdf/PR/v95/i6/p1586_1.

Dunson:2020:HAF

[DJ20] D. B. Dunson and J. E. Johndrow. The Hastings algorithm at fifty. Biometrika, 107(1):1–23, March 2020. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://academic.oup.com/biomet/article-abstract/107/1/
1/5686745.

Demuth:1953:M

[DJK^{+}53] Howard B. Demuth, John B. Jackson, Edmund Klein, N. Metropolis, Walter Orvedahl, and James H. Richardson. MANIAC. In Proceedings of the Association for Computing Machinery, Toronto, 1952, pages 13–17. Sauls Lithograph Co. (for the Association for Computing Machinery), Washington, DC, USA, 1953.

Diaconis:2004:NRM

[DN04] Persi Diaconis and J. W. Neuberger. Numerical results for the Metropolis algorithm. Experimental Mathematics, 13(2):207–214, ???? 2004. CODEN ???? ISSN 1058-6458 (print), 1944-

950X (electronic). URL http://projecteuclid.org/euclid.em/
1090350935.

Diaconis:1995:WDW

[DSC95] Persi Diaconis and Laurent Saloff-Coste. What do we know about the Metropolis algorithm? In ACM [ACM95], pages 112–129. ISBN 0-89791-718-9. LCCN QA 76.6 A13 1995.

URL http://www.acm.org/pubs/articles/proceedings/stoc/
225058/p112-diaconis/p112-diaconis.pdf;   http://www.acm.
org/pubs/citations/proceedings/stoc/225058/p112-diaconis/

. ACM order no. 508950.

Diaconis:1998:WDW

[DSC98] P. Diaconis and L. Saloff-Coste. What do we know about the Metropolis Algorithm? Journal of Computer and System Sciences, 57(1):20–36, August 1998. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022000098915768.

Dyson:2012:MC

[Dys12a] George Dyson. Monte Carlo. In Turing’s cathedral: the origins of the digital universe [Dys12b], chapter 10, pages 175–199. ISBN 0-375-42277-3 (hardcover). LCCN QA76.17 .D97 2012. Pages 191– 192 describe the origin of the Monte Carlo method.

Dyson:2012:TCO

[Dys12b] George Dyson. Turing’s cathedral: the origins of the digital universe. Pantheon Books, New York, NY, USA, 2012. ISBN 0-375-42277-3 (hardcover). xxii + 401 pp. LCCN QA76.17 .D97 2012.

Dyson:2012:UD

[Dys12c] George Dyson. Ulam’s demons. In Turing’s cathedral: the origins of the digital universe [Dys12b], chapter 11, pages 200–224. ISBN 0-375-42277-3 (hardcover). LCCN QA76.17 .D97 2012. Pages 191– 192 describe the origin of the Monte Carlo method.

Edwards:1985:BRS

[Edw85] John S. Edwards. Book reviews: Software Lifecycle Management. The Incremental Method, by William C. Cave and Gilbert W. Maymon, Science, Computers and the Information Onslaught, by Donald M. Kerr, Karl Braithwaite, N. Metropolis, David H. Sharp, and Gian-Carlo Rota, My Personal Computer and Other Family Crises, by Ben Ross Schneider, Jr. OR: the journal of the Operational Research Society, 36(11):1075–1076, November 1985.

CODEN OPRQAK. ISSN 0160-5682 (print), 1476-9360 (elec-

tronic). URL http://links.jstor.org/sici?sici=0160-5682%
28198511%2936%3A11%3C1075%3ASLMTIM%3E2.0.CO%3B2-S. See

[KBM^{+}84].

Everett:1971:ATR

[EM71a] C. J. Everett and N. Metropolis. Approximation of the νth root of N. Studies in Applied Mathematics, 50:189–191, 1971. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic).

Everett:1971:GGL

[EM71b] C. J. Everett and N. Metropolis. A generalization of the Gauss limit for iterated means. Advances in Mathematics, 7:297–300, 1971. CODEN ADMTA4. ISSN 0001-8708 (print), 1090-2082 (electronic).

Everett:1972:CNM

[EM72] C. J. Everett and N. Metropolis. On completely normal (0, 1)- matrices and symmetrizability. Journal of Combinatorial Theory (Series A), 13:367–373, 1972. CODEN JCBTA7. ISSN 0097-3165 (print), 1096-0899 (electronic).

Everett:1975:RXP

[EM75] C. J. Everett and N. Metropolis. On the roots of X_{m} + 1 in the p-adic field Q_{p}. Notices of the American Mathematical Society, 22(6):A619, October 1975. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Fano:1960:BRJ

[Fan60] U. Fano. Book review: The 3-j and 6-j Symbols, by Manuel Rotenberg; R. Bivins; N. Metropolis; John K. Wooten, Jr. Science (New Series), 132(3420):143, July 1960. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://links.jstor.org/

sici?sici=0036-8075%2819600715%293%3A132%3A3420%3C143%
3AT3A6S%3E2.0.CO%3B2-F. See [RBMW59].

Frankel:1947:CLD

[FM47] S. Frankel and N. Metropolis. Calculations in the liquid-drop model of fission. Physical Review, Series 2, 72(10):914–925, November 15, 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Fermi:1952:NSM

[FM52] Enrico Fermi and Nicholas Metropolis. Numerical solution of a minimum problem. Report LA-1492, US Atomic Energy Commission, Washington, DC, USA, November 19, 1952.

Fraser:1968:AUA

[FM68] M. Fraser and N. Metropolis. Algorithms in unnormalized arithmetic. III. Matrix inversion. Numerische Mathematik, 12(5):416– 428, December 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Fermi:1954:PSA

[FMA54] E. Fermi, N. Metropolis, and E. F. Alei. Phase shift analysis of the scattering of negative pions by hydrogen. Physical Review, 95(6):1581–1585, September 15, 1954. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/pdf/PR/v95/i6/p1581_1.

Faltin:1975:RNW

[FMRR75] F. Faltin, N. Metropolis, B. Ross, and G.-C. Rota. The real numbers as a wreath product. Advances in Mathematics, 16(3):278– 304, 1975. CODEN ADMTA4. ISSN 0001-8708 (print), 1090-2082 (electronic).

Faltin:1976:RNW

[FMRR76] F. Faltin, N. Metropolis, B. Ross, and G.-C. Rota. The real numbers as a wreath product. In Metropolis et al. [MOR76], pages 271–297. ISBN 0-12-492150-7. LCCN QA1 .L588 1974.

Faltin:1976:BAA

[FMRR77a] F. Faltin, N. Metropolis, B. Ross, and G.-C. Rota. A Boolean analysis of addition and multiplication. Studies in Applied Mathematics, 56(2):147–158, 1976/1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic). URL http:/

/www.osti.gov/energycitations/product.biblio.jsp?osti_
id=6234187&query_id=0.

Faltin:1977:BAA

[FMRR77b] F. Faltin, N. Metropolis, B. Ross, and G.-C. Rota. Boolean analysis of addition and multiplication. Studies in Applied Mathematics, 56(2):147–158, ???? 1977. CODEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic).

Feynman:1949:ESEa

[FMT49a] R. P. Feynman, N. Metropolis, and E. Teller. Equations of state of elements based on the generalized Fermi–Thomas theory. Technical Report AECD-2448, Technical Information Branch, Oak Ridge Operations, AEC, Oak Ridge, TN, USA, January 20,

1949. 41 pp. URL http://www.osti.gov/accomplishments/
documents/fullText/ACC0107.pdf;   http://www.osti.gov/
servlets/purl/4417654-BCgOtj/native/.

Feynman:1949:ESEb

[FMT49b] R. P. Feynman, N. Metropolis, and E. Teller. Equations of state of elements based on the generalized Fermi–Thomas theory. Physical Review, Series 2, 75(10):1561–1573, May 15, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Frankel:1950:IDT

[FMT50] Stanley Frankel, Nicholas Metropolis, and Anthony Turkevich. Ignition of deuterium-tritium mixtures: Numerical calculations using the ENIAC. Classified report LA-525, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, March 2, 1950. 23 pp.

Fartaria:2006:TSA

[FNR^{+}06] Rui P. S. Fartaria, Rodrigo S. Neves, Pedro C. R. Rodrigues, Filomena F. M. Freitas, and Fernando M. S. Silva Fernandes. A time saving algorithm for the Monte Carlo method of Metropolis. Computer Physics Communications, 175(2):116–121, July 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506001147. See erratum [FNR^{+}07].

Fartaria:2007:ETS

[FNR^{+}07] Rui P. S. Fartaria, Rodrigo S. Neves, Pedro C. R. Rodrigues, Filomena F. M. Freitas, and Fernando M. S. Silva Fernandes. Erratum to “A time saving algorithm for the Monte Carlo method of Metropolis” [Computer Physics Communications 175 (2006) 116– 121]. Computer Physics Communications, 176(3):250, February 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506004176. See [FNR^{+}06].

Gardiner:1956:CSI

[GLMU56] Verna Gardiner, R. Lazarus, N. Metropolis, and S. Ulam. On certain sequences of integers defined by sieves. Mathematics Magazine, 29:117–122, 1956. CODEN MAMGA8. ISSN 0011-

801x. URL http://links.jstor.org/sici?sici=0025-570X%
28195601%2F02%2929%3A3%3C117%3AOCSOID%3E2.0.CO%3B2-H.

Gamow:1954:NPC

[GM54] G. Gamow and N. Metropolis. Numerology of polypeptide chains. Science (New Series), 120(3124):779–780, November 12, 1954. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

jstor.org/stable/1681634.

Gardiner:1970:CAC

[GM70a] V. Gardiner and N. Metropolis. A comprehensive approach to computer arithmetic. Technical Report LA-4531, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1970.

Gardiner:1970:SDA

[GM70b] V. Gardiner and N. Metropolis. Significant digit arithmetic on a CDC 6600. Technical Report LA-4470, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, January 1, 1970.

4 pp. URL http://www.osti.gov/energycitations/product.
biblio.jsp?osti_id=4093265&query_id=0.

Gubernatis:1986:PCF

[GM86] James E. Gubernatis and N. Metropolis, editors. Proceedings of the Conference on Frontiers of Quantum Monte Carlo, Los Alamos National Laboratory, Los Alamos, New Mexico, September 3–6, 1985, volume 43(5/6) of Journal of statistical physics,; v. Plenum Press, New York, NY, USA; London, UK, 1986. ISSN 0022-4715 (print), 1572-9613 (electronic).

Greenwood:1955:CCT

[Gre55] Robert E. Greenwood. Coupon collector’s test for random digits. Mathematical Tables and Other Aids to Computation, 9(49):1–5, January 1955. CODEN MTTCAS. ISSN 0891-6837. URL http:/

/www.jstor.org/stable/2002211.

Greene:1984:CA

[Gre84] Curtis Greene, editor. Combinatorics and algebra: Proceedings of the AMS-NSF Joint Summer Research Conference on Combinatorics and Algebra, held at the University of Colorado, Boulder, during June 5–11, 1983, volume 34 of Contemporary mathematics. American Mathematical Society, Providence, RI, USA, 1984. ISBN 0-8218-5029-6 (paperback). ISSN 0271-4132 (print), 1098-3627 (electronic). LCCN QA150 .C647 1984.

Gubernatis:2005:MRM

[Gub05] J. E. Gubernatis. Marshall Rosenbluth and the Metropolis algorithm. Physics of Plasmas, 057303:5, May 2005. CODEN PH-PAEN. ISSN 1070-664X (print), 1089-7674 (electronic), 1527-2419.

Gustafson:1998:GWM

[Gus98] Paul Gustafson. A guided walk Metropolis algorithm. Statistics and Computing, 8(4):357–364, December 1998. CODEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1008880707168.

Hammer:1965:BRBa

[Ham65a] Preston C. Hammer. Book review: Experimental Arithmetic, High Speed Computing and Mathematics by N. C. Metropolis; A. H. Taub; John Todd; C. B. Tompkins. Technometrics, 7(1):82, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266139.

Hammer:1965:BRE

[Ham65b] Preston C. Hammer. Book review: Experimental Arithmetic, High Speed Computing and Mathematics by N. C. Metropolis, A. H. Taub, John Todd, and C. B. Tompkins. Technometrics, 7(1):82, February 1965. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://links.jstor.org/sici?sici=
0040-1706%28196502%297%3A1%3C82%3AEAHSCA%3E2.0.CO%3B2-
9.

Hargittai:2010:JET

[Har10] István Hargittai. Judging Edward Teller: a closer look at one of the most influential scientists of the Twentieth Century. Prometheus Books, Amherst, NY, USA, 2010. ISBN 1-61614-221-9 (hardcover). 575 pp. LCCN QC16.T37 H37 2010. Afterword by Richard Garwin.

Hastings:1970:MCS

[Has70] W. K. Hastings. Monte Carlo sampling methods using Markov chains and their applications. Biometrika, 57(1):97–109, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334940;

http://www.probability.ca/hastings/. This paper introduces

what is now known as the Metropolis–Hastings algorithm, a generalization of the work in [MRR^{+}53]. See [Hit03, page 255], and [DJ20] for a 50-year review.

Hyman:1996:DAM

[HBLM96] J. Hyman, W. Beyer, J. Louck, and N. Metropolis. Development of the applied mathematics originating from the group theory of physical and mathematical problems. Technical report LA-UR–96-1880, Los Alamos National Laboratory, Los Alamos, NM,

USA, 1996. 7 pp. URL http://www.osti.gov/energycitations/
product.biblio.jsp?osti_id=257450&query_id=0;   http:/
/www.osti.gov/servlets/purl/257450-biVIAW/webviewable/.

Heinzmann:1999:ONM

[Hei99] David Heinzmann. Obituary: Nick Metropolis. Bulletin — Institute of Mathematical Statistics, 28(6):??, November/December 23, 1999. CODEN SMBCVA. ISSN 0146-3942. URL http://www.

imstat.org/bulletin/dec1999/node5.html.

Hitchcock:2003:HMH

[Hit03] David B. Hitchcock. A history of the Metropolis–Hastings algorithm. The American Statistician, 57(4):254–257, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/asa/00031305/v57n4/s7/p254;
http://www.jstor.org/stable/30037292.

Harlow:1983:CCW

[HM83] Francis H. Harlow and N. Metropolis. Computing & computers: Weapons simulation leads to the computer era. Los Alamos Science, 4(7):132–141, Winter/Spring 1983. CODEN LASCDI. ISSN

0273-7116. URL http://library.lanl.gov/cgi-bin/getfile?
00285876.pdf;   http://library.lanl.gov/cgi-bin/getfile?
number7.htm.

Hoffman:1955:STC

[HMG55] Joseph G. Hoffman, Nicholas Metropolis, and Verna Gardiner. Study of tumor cell populations by Monte Carlo methods. Science (New Series), 122(3167):465–466, September 9, 1955. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.

org/stable/1750182.

Hoffman:1956:DCS

[HMG56] J. G. Hoffman, N. Metropolis, and V. Gardiner. Digital computer studies of cell multiplication by Monte Carlo methods. Journal of the National Cancer Institute: JNCI, 17(2):175–188, August 1, 1956. CODEN JNCIAM. ISSN 0027-8874 (print), 1460-2105

(electronic). URL http://www.osti.gov/energycitations/
product.biblio.jsp?osti_id=4350852&query_id=0.

Herbst:1955:APC

[HMW55] Eugene H. Herbst, N. Metropolis, and Mark B. Wells. Analysis of problem codes on the MANIAC. Mathematical Tables and Other Aids to Computation, 9(49):14–20, January 1955. CODEN MTTCAS. ISSN 0891-6837. URL http://

links.jstor.org/sici?sici=0891-6837%28195501%299%3A49%
3C14%3AAOPCOT%3E2.0.CO%3B2-C.

Holden:1998:GCM

[Hol98] Lars Holden. Geometric convergence of the Metropolis–Hastings simulation algorithm. Statistics & Probability Letters, 39(4):371– 377, August 21, 1998. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715298000960.

Haigh:2014:ABE

[HPR14] T. Haigh, M. Priestley, and C. Rope. Los Alamos bets on ENIAC: Nuclear Monte Carlo simulations, 1947–1948. IEEE Annals of the History of Computing, 36(3):42–63, July 2014. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hecker:2000:EFH

[HR00] Siegfried S. Hecker and Gian-Carlo Rota, editors. Essays on the Future: in honor of Nick Metropolis. Birkhäuser Boston Inc., Cambridge, MA, USA, 2000. ISBN 1-4612-0777-0 (e-book), 1-4612-6898-2 (print), 0-8176-3856-3, 3-7643-3856-3. xvii + 276 pp. LCCN Q160.2 .E77 2000.

Haario:2001:AMA

[HST01] Heikki Haario, Eero Saksman, and Johanna Tamminen. An adaptive Metropolis algorithm. Bernoulli, 7(2):223–242, April 2001. CODEN ???? ISSN 1350-7265 (print), 1573-9759 (electronic).

URL http://projecteuclid.org/euclid.bj/1080222083.

Hurd:1985:NEM

[Hur85] Cuthbert C. Hurd. A note on early Monte Carlo computations and scientific meetings. Annals of the History of Computing, 7(2):141–155, April/June 1985. CODEN AH-

COE5. ISSN 0164-1239. URL http://dlib.computer.org/an/
books/an1985/pdf/a2141.pdf;   http://www.computer.org/
annals/an1985/a2141abs.htm. Includes typeset reprint of

[RUvN47].

IEEE:1972:ITS

[IEE72a] IEEE, editor. 2nd IEEE-TCCA Symposium on Computer Arithmetic, College Park, Maryland, May 15–16, Maryland. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1972. ISBN ???? LCCN ????

IEEE:1972:IAD

[IEE72b] IEEE, editor. Innovative architecture: digest of papers: COMP-CON 72, 6. annual IEEE Computer Society International Conference, Jack Tar Hotel, San Francisco, California, September 12–14, 1972. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1972. LCCN TK7885.A1 C53 1972. IEEE order number 72CH0659-3C.

IEEE:1975:SCA

[IEE75] 3rd Symposium on Computer Arithmetic, November 19–20, 1975, Southern Methodist University, Dallas, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1975. LCCN QA76.6.S919 1975. IEEE order number CH1017-3C.

Metropolis:1959:ITC

[ITC74] Nicholas Constantine Metropolis, E. Priory, and S. Ulam, editors. International tracts in computer science and technology and their application, 1959–1974. ISSN 0074-9141. Pergamon Press, New York, NY, USA.

Jacobs:1977:SAN

[Jac77] David A. H. Jacobs, editor. The State of the Art in Numerical Analysis: Proceedings of the Conference on the State of the Art in Numerical Analysis held at the University of York, April 12th-15th, 1976. Academic Press, New York, NY, USA, 1977. ISBN 0-12-378650-9. LCCN QA297 .C646 1976.

Jackson:1951:M

[JM51] John B. Jackson and N. Metropolis. The MANIAC. Report LA-1725, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, December 15, 1951. ii + 306 pp. URL http://www.bitsavers.

org/pdf/lanl/LA-1725_The_MANIAC_Jul54.pdf.

Jackson:1954:M

[JM54] John B. Jackson and N. Metropolis. The MANIAC. Technical report, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1954. 306 pp. Report LA-1725.

Juncosa:1964:BRB

[Jun64a] Mario Juncosa. Book review: Proceedings of the 15th Symposium in Applied Mathematics of the A.M.S. Experimental Arithmetic, High Speed Computing and Mathematics (N. C. Metropolis, A. H. Taub, John Todd C. B. Tompkins). SIAM Review, 6(4):468–471, ???? 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Juncosa:1964:BRP

[Jun64b] Mario Juncosa. Book review: Proceedings of the 15th Symposium in Applied Mathematics of the A.M.S. Experimental Arithmetic, High Speed Computing and Mathematics, by N. C. Metropolis, A. H. Taub, John Todd, and C. B. Tompkins. SIAM Review, 6(4):468–471, October 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:/

/links.jstor.org/sici?sici=0036-1445%28196410%296%3A4%
3C468%3APOT1SI%3E2.0.CO%3B2-T.

Kalenich:1965:IPP

[Kal65] Wayne A. Kalenich, editor. Information processing 1965: proceedings of IFIP congress 65; New York City May 24–29, 1965. Spartan Books, Washington, DC, USA, 1965. LCCN ???? Two volumes.

Kerr:1984:SCI

[KBM^{+}84] Donald M. Kerr, Karl Braithwaite, N. Metropolis, David H. Sharp, and Gian-Carlo Rota, editors. Science, computers, and the information onslaught: a collection of essays. Academic Press, New York, NY, USA, 1984. ISBN 0-12-404970-2. xiii + 276 pp. LCCN Q223 .S24 1984.

Kilpatrick:1954:SVC

[KKHM54] John E. Kilpatrick, William E. Keller, Edward F. Hammel, and Nicholas Metropolis. Second virial coefficients of He^{3} and He^{4}. Physical Review, Series 2, 94(5):1103–1110, June 1, 1954. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/pdf/PR/v94/i5/p1103_1.

Konopinski:1943:CAU

[KMTW43] E. Konopinski, N. Metropolis, E. Teller, and L. Woods. Critical amounts of uranium compounds. Technical Report CF-548, United States Atomic Energy Commission, Washington, DC, USA, March 19, 1943. 13 pp. URL http://www.osti.gov/

accomplishments/pdf/DE04384835/DE04384835.pdf;   http:
//www.osti.gov/servlets/purl/4384835-NQvnzm/native/.

Lee:1981:RMA

[LB81] J. A. N. Lee and Hal Berghel. Reviews: N. Metropolis, et al.: A History of Computing in The Twentieth Century; capsule reviews. Annals of the History of Computing, 3(2): 193–208, April/June 1981. CODEN AHCOE5. ISSN 0164-

1239. URL http://dlib.computer.org/an/books/an1981/pdf/
a2193.pdf;   http://www.computer.org/annals/an1981/a2193abs.
htm.

Liang:2010:DMH

[Lia10] Faming Liang. A double Metropolis–Hastings sampler for spatial models with intractable normalizing constants. Journal of Statistical Computation and Simulation, 80(9):1007–1022, 2010. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Lide:2001:CEM

[Lid01] D. R. Lide, editor. A Century of Excellence in Measurements, Standards, and Technology: A Chronicle of Selected NBS/NIST Publications, 1901–2000, volume 958. National Technical Information Service, Washington, DC, USA, 2001. ix + 386

pp. URL https://nvlpubs.nist.gov/nistpubs/sp958-lide/

cntsp958old.htm. NIST Special Publication.

Louck:1980:HSN

[LM80] J. D. Louck and N. Metropolis. Hidden symmetry and the numbertheoretic structure of the energy levels of a perturbed harmonic oscillator. Advances in Applied Mathematics, 1(2):182–220, 1980. ISSN 0196-8858 (print), 1090-2074 (electronic).

Louck:1981:NTD

[LM81] J. D. Louck and N. Metropolis. Number-theoretic degeneracy of the energy levels of a perturbed N-dimensional isotropic harmonic oscillator. Advances in Applied Mathematics, 2(2):138–171, 1981. ISSN 0196-8858 (print), 1090-2074 (electronic). URL http:/

/www.osti.gov/energycitations/product.biblio.jsp?osti_
id=6233895&query_id=0.

Louck:1986:CDL

[LM86a] J. D. Louck and N. Metropolis. Classical dynamics and Lie groups. American Journal of Physics, 54(6):558–564, June 1986. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://ajp.aapt.org/resource/1/ajpias/v54/i6/p558_s1.

Louck:1986:SDT

[LM86b] James D. Louck and N. (Nicholas) Metropolis. Symbolic Dynamics of Trapezoidal Maps, volume 27 of Mathematics and its Applications. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1986. ISBN 90-277-2197-1. viii + 312 pp. LCCN QA331 .L813 1986.

Lazarus:1956:MI

[LMO^{+}56] R. B. Lazarus, N. Metropolis, W. Orvedahl, J. H. Richardson, W. Spack, Jr., R. L. Bivins, J. V. Caulfield, I. Kral, A. F. Malmberg, G. T. McKinley, and R. E. Williamson. MANIAC II. Report LA-2083, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, October 1, 1956. 54 pp. URL https://sgp.fas.org/

othergov/doe/lanl/lib-www/la-pubs/00320765.pdf.

Ledin:1981:RJG

[LTC^{+}81] George Ledin, Jr., Henry S. Tropp, Paul Ceruzzi, Martin Campbell-Kelly, and K. W. Smillie. Reviews: J. G. Santesmases: Obra e Inventos de Torres Quevedo; C. Evans: Pioneers of Computing; N. Metropolis, et al.: A History of Computing in the Twentieth Century; capsule reviews. Annals of the History of Computing, 3(4):416–430, October/December 1981. CODEN AHCOE5.

ISSN 0164-1239. URL http://dlib.computer.org/an/books/
an1981/pdf/a4416.pdf.

Metropolis:1958:SDC

[MA58] N. Metropolis and R. L. Ashenhurst. Significant digit computer arithmetic. IRE Transactions on Electronic Computers, EC-7 (4):265–267, December 1958. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
tp=&arnumber=5222657.

Metropolis:1963:BOU

[MA63] N. Metropolis and R. L. Ashenhurst. Basic operations in an unnormalized arithmetic system. IEEE Transactions on Electronic Computers, EC-12(6):896–904, December 1963. CODEN

IEECA8. ISSN 0367-7508. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4038037.

Metropolis:1965:RCU

[MA65] N. Metropolis and R. L. Ashenhurst. Radix conversion in an unnormalized arithmetic system. Mathematics of Computation, 19 (91):435–441, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://links.jstor.org/

sici?sici=0025-5718%28196507%2919%3A91%3C435%3ARCIAUA%
3E2.0.CO%3B2-D.

Metropolis:1987:OHI

[MA87] N. Metropolis and William Aspray. Oral history interview with Nicholas Metropolis. Audio recording, 1987. The Charles Babbage Institute.

MacKenzie:1991:IAL

[Mac91] Donald MacKenzie. The influence of the Los Alamos and Livermore National Laboratories on the development of supercomputing. Annals of the History of Computing, 13(2): 179–201, April/June 1991. CODEN AHCOE5. ISSN 0164-

1239. URL http://dlib.computer.org/an/books/an1991/pdf/
a2179.pdf;   http://www.computer.org/annals/an1991/a2179abs.
htm.

Marshall:1956:INM

[Mar56] Andrew W. Marshall. An introductory note [on Monte Carlo method]. In Meyer [Mey56], pages 14–?? LCCN QA273 U577.

May:1976:SMM

[May76] Robert M. May. Simple mathematical models with very complicated dynamics. Nature, 261(5560):459–467, June 10, 1976.

CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v261/
n5560/pdf/261459a0.pdf.

Metropolis:1958:MCCb

[MBS^{+}58a] N. Metropolis, R. Bivins, M. Storm, J. M. Miller, G. Friedlander, and Anthony Turkevich. Monte Carlo calculations on intranuclear cascades. II. High-energy studies and pion processes. Physical Review, 110(1):204–219, April 1, 1958. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/pdf/PR/v110/i1/p204_1.

Metropolis:1958:MCC

[MBS^{+}58b] N. Metropolis, R. Bivins, M. Storm, Anthony Turkevich, J. M. Miller, and G. Friedlander. Monte Carlo calculations on intranuclear cascades. I. Low-energy studies. Physical Review, 110(1):185– 203, April 1, 1958. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/pdf/PR/
v110/i1/p185_1;   http://www.osti.gov/energycitations/
product.biblio.jsp?osti_id=4331465&query_id=0.

Metropolis:1939:NUB

[Met39] Nicholas Metropolis. A new ultraviolet band system of silver iodide. Physical Review, 55(7):636–638, April 1, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/pdf/PR/v55/i7/p636_1.

Metropolis:1941:SEBa

[Met41a] N. Metropolis. I. The structure of electronic bands of polyatomic molecules: II. Vibrational analysis of the absorption system of sulphur dioxide of λ3400–2600. Ph.D. thesis, Department of Physics, University of Chicago, Chicago, IL, USA, 1941. 301 pp. Thesis advisor, Robert S. Mulliken.

Metropolis:1941:SEBb

[Met41b] N. Metropolis. The structure of electronic bands of polyatomic molecules. I. prolate approximation for XY_{2} molecules. Physical Review, 60(4):283–294, August 15, 1941. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/pdf/PR/v60/i4/p283_1.

Metropolis:1941:VAA

[Met41c] N. Metropolis. Vibrational analysis of the absorption system of sulphur dioxide at λ3400–2600. Physical Review, 60(4):295–301, August 15, 1941. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/pdf/PR/v60/i4/
p295_1.

Metropolis:1950:BRG

[Met50] N. Metropolis. Book review: Giant Brains or Machines That Think, by Edmund C. Berkeley. Journal of the American Statistical Association, 45(252):573–574, December 1950. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL

http://www.jstor.org/stable/2280729.

Metropolis:1954:EFO

[Met54] N. Metropolis. Enrico Fermi offprints (1922–1954). Technical report, University of Chicago, Chicago, IL, USA, 1954. From the papers of Herbert Anderson.

Metropolis:1955:EF

[Met55] Nicholas Metropolis. Enrico Fermi. Physics Today, 8(11):10–12, November 1955. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://www.physicstoday.org/
resource/1/phtoad/v8/i11/p10_s1.

Metropolis:1956:PSM

[Met56] N. Metropolis. Phase shifts — middle squares — wave equations. In Meyer [Mey56], pages 29–36. LCCN QA273 U577.

Metropolis:1960:SLO

[Met60] N. C. Metropolis. Symposium on the logical organization of very high speed computers. In Information processing: proceedings of the International Conference on Information Processing, Paris, France, 15–20 June 1959, pages 432–436. UNESCO, Paris, France, 1960. LCCN QA76 .I578.

Metropolis:1964:IQRb

[Met64a] N. C. Metropolis. ICR quarterly report. Technical Report TID-21413, Institute for Computer Research, University of Chicago, Chicago, IL, USA, August 1, 1964. 231 pp. URL http:/

/www.osti.gov/energycitations/product.biblio.jsp?osti_
id=4659841&query_id=0.

Metropolis:1964:IQRa

[Met64b] N. C. Metropolis. ICR quarterly report, number 1. Technical Report TID-21595, Institute for Computer Research, University of Chicago, Chicago, IL, USA, May 1, 1964. 136

pp. URL http://www.osti.gov/energycitations/product.
biblio.jsp?osti_id=4656875&query_id=0.

Metropolis:1965:AUA

[Met65a] N. Metropolis. Algorithms in unnormalized arithmetic. I: Recurrence relations. Numerische Mathematik, 7(2):104–112, April 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). See erratum [Met65c].

Metropolis:1965:AIE

[Met65b] N. Metropolis. Analysis of inherent errors in matrix decomposition using unnormalized arithmetic. In Kalenich [Kal65], pages 441– 442. LCCN ???? Two volumes.

Metropolis:1965:BAU

[Met65c] N. Metropolis. Berichtigung: Algorithms in unnormalized arithmetic. I: Recurrence relations. Numerische Mathematik, 7(4):354, August 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Metropolis:1965:RRB

[Met65d] N. Metropolis. Review: R65-35 Binary-Compatible Signed-Digit Arithmetic by Algirdas Avisienis. IEEE Transactions on Electronic Computers, EC-14(3):499, June 1965. CODEN IEECA8. ISSN

0367-7508. URL http://ieeexplore.ieee.org/iel5/4037753/
4038444/04038481.pdf.

Metropolis:1965:IQR

[Met65e] N. C. Metropolis. ICR quarterly report, number 4. Technical Report COO-614-31, Institute for Computer Research, University of Chicago, Chicago, IL, USA, February 1, 1965. 89

pp. URL http://www.osti.gov/energycitations/product.
biblio.jsp?osti_id=4632649&query_id=0.

Metropolis:1968:ANA

[Met68] N. Metropolis. Algorithms in un-normalized arithmetic: Polynomial evaluation and matrix decomposition. Colloques internationaux, Centre National de la Recherche Scientifique, Paris, 165: 293–303, 1968.

Metropolis:1972:ABCa

[Met72a] N. Metropolis. Analyzed binary computing. Technical Report LA-DC–72-783; CONF-720916–2, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, September 12, 1972. 4

pp. URL http://www.osti.gov/energycitations/product.
biblio.jsp?osti_id=4647144&query_id=0.

Metropolis:1972:ABCb

[Met72b] N. Metropolis. Analyzed binary computing. In IEEE [IEE72b], pages 81–84. LCCN TK7885.A1 C53 1972. IEEE order number 72CH0659-3C.

Metropolis:1972:ABCc

[Met72c] Nicholas C. Metropolis. Analyzed binary computing. In IEEE [IEE72a], pages 1–14. ISBN ???? LCCN ???? URL http://

ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=6153912.

Metropolis:1973:ABC

[Met73] Nicholas C. Metropolis. Analyzed binary computing. IEEE Transactions on Computers, C-22(6):573–576, June 1973. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
tp=&arnumber=5009109.

Metropolis:1976:BRA

[Met76a] N. Metropolis. Book review: Adventures of a Mathematician, by S. M. Ulam. Science (New Series), 193(4254):568– 569, August 1976. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://links.jstor.org/sici?sici=0036-8075%
2819760813%293%3A193%3A4253%3C568%3AM%3E2.0.CO%3B2-M.

Metropolis:1976:M

[Met76b] N. Metropolis. Memoirs. Science (New Series), 193(4253):568–569, ???? 1976. ISSN 0036-8075 (print), 1095-9203 (electronic).

Metropolis:1976:MSA

[Met76c] N. Metropolis. Methods of significance arithmetic. Technical report LA-UR-76-661;CONF-760428-1, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, January 1, 1976. URL http:/

/www.osti.gov/energycitations/product.biblio.jsp?osti_

id=7189580&query_id=0. Presented at the Conference on the state of the art in numerical analysis, 12 April 1976, University of York, England, UK.

Metropolis:1977:MSA

[Met77] N. Metropolis. Methods of significance arithmetic. In Jacobs [Jac77], pages 179–192. ISBN 0-12-378650-9. LCCN QA297 .C646 1976.

Metropolis:1978:MGB

[Met78] N. C. Metropolis. Maniac — great big toy. Datamation, 24(8):80, ???? 1978. CODEN DTMNAT. ISSN 0011-6963.

Metropolis:1979:SIN

[Met79] N. Metropolis. Summation of imprecise numbers. Technical report, U.S. Department of Energy, Washington, DC, USA, July 1979. 5 pp. LA-7916-MS.

Metropolis:1980:M

[Met80a] N. Metropolis. The MANIAC. In Metropolis et al. [MHR80], pages 457–464. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Metropolis:1980:SIN

[Met80b] N. Metropolis. Summation of imprecise numbers. Computers and Mathematics with Applications, 6(3):297–299, 1980. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Metropolis:1985:MCB

[Met85] Nicholas Metropolis. Monte Carlo: in the beginning and some great expectations. In Alcouffe et al. [A^{+}85], pages 62–70. ISBN 0-387-16070-1 (paperback). LCCN QC20.7.M65 M65 1985. DM66.00 (West Germany).

Metropolis:1987:BMC

[Met87a] N. Metropolis. The beginning of the Monte Carlo method. Los Alamos Science, 15 (Special Issue, Stanisław Ulam 1909–1984): 125–130, 1987. CODEN LASCDI. ISSN 0273-7116. URL http://

library.lanl.gov/cgi-bin/getfile?00326866.pdf;   http:
//library.lanl.gov/cgi-bin/getfile?15-12.pdf;   http:/
/library.lanl.gov/la-pubs/00326866.pdf.

Metropolis:1987:AE

[Met87b] N. Metropolis. The Los Alamos experience, 1943–1954. Technical report LA-UR-87-1353, Los Alamos National Laboratory, Los Alamos, NM, USA, May 1987. 18 pp. Paper presented to the ACM Conference on History of Scientific and Numeric Computation, Los Alamos National Laboratory, Los Alamos, NM, USA.

Metropolis:1990:AE

[Met90] N. Metropolis. The Los Alamos experience, 1943–1954. In Nash [Nas90], pages 237–250. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Metropolis:1992:ACP

[Met92] N. (Nicholas) Metropolis. The age of computing: a personal memoir. Dædalus, 121(1):119–130, Winter 1992. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://www.

bio.net/hypermail/bioforum/1992-September/000238.html;
 http://www.jstor.org/stable/20025423.

Metropolis:1993:ACP

[Met93] N. Metropolis. The age of computing: a personal memoir. In Metropolis and Rota [MR93], pages 119–130. ISBN 0-262-63154-7. LCCN QA76.58 .N48 1993. US$13.95.

Metropolis:1996:RR

[Met96] Nicholas Metropolis. Random reminiscences. In Behind tall fences: stories and experiences about Los Alamos at its beginning [Ano96], pages 69–86. ISBN 0-941232-91-3. LCCN QC773.3.U5 B43 1996.

Meyer:1956:SMC

[Mey56] Herbert A. Meyer, editor. Symposium on Monte Carlo Methods: held at the University of Florida, March 16 and 17, 1954. Wiley, New York, NY, USA, 1956. LCCN QA273 U577.

Metropolis:1980:HCT

[MHR80] Nicholas Metropolis, Jack Howlett, and Gian-Carlo Rota, editors. A History of Computing in the Twentieth Century: a Collection of Essays. Academic Press, New York, NY, USA, 1980. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Metropolis:1987:NDP

[MKR87] N. (Nicholas) Metropolis, Donald M. Kerr, and Gian-Carlo Rota, editors. New Directions in Physics: The Los Alamos 40th Anniversary Volume. Academic Press, New York, NY, USA, 1987. ISBN 0-12-492155-8. xii + 292 pp. LCCN QC44 .N49

1987. URL http://www.osti.gov/energycitations/product.
biblio.jsp?osti_id=6120718&query_id=0.

Menzel:1967:AUA

[MM67] M. Menzel and N. Metropolis. Algorithms in unnormalized arithmetic. II. Unrestricted polynomial evaluation. Numerische Mathematik, 10(5):451–462, November 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Metropolis:1982:ECA

[MN82] N. Metropolis and E. C. Nelson. Early computing at Los Alamos. Annals of the History of Computing, 4(4):348–357, October/December 1982. CODEN AHCOE5. ISSN 0164-

1239. URL http://dlib.computer.org/an/books/an1982/pdf/
a4348.pdf;   http://ieeexplore.ieee.org/stamp/stamp.jsp?
tp=&arnumber=4640758;   http://www.computer.org/annals/
an1982/a4348abs.htm.

Metropolis:1981:NCS

[MNR81] N. Metropolis, G. Nicoletti, and Gian-Carlo Rota. A new class of symmetric functions. In Nachbin [Nac81], pages 563–575. ISBN 0-12-512801-0 (vol. 1), 0-12-512802-9 (vol. 2). LCCN A300 .M294.

Metropolis:1976:SAM

[MOR76] N. (Nicholas) Metropolis, Steven A. Orszag, and Gian-Carlo Rota, editors. Surveys in applied mathematics: Essays dedicated to S. M. Ulam: proceedings of the first Los Alamos Symposium on Mathematics in the Natural Sciences, Surveys in applied mathematics. Academic Press, New York, NY, USA, 1976. ISBN 0-12-492150-7. LCCN QA1 .L588 1974.

Metropolis:1966:SSC

[MPF66] N. Metropolis, J. R. Pasta, and M. Fraser. Simulation of the SEL-810A computer on MANIAC II (Selma). Technical Report LA-3640, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, January 1, 1966. 38 pp. URL http:/

/www.osti.gov/energycitations/product.biblio.jsp?osti_
id=4440527&query_id=0.

Metropolis:1950:TAM

[MR50] N. Metropolis and George W. Reitwiesner. Table of atomic masses. Technical Report USAEC NP-1980, Argonne National Laboratory, Argonne, IL, USA, March 1, 1950. 472

pp. URL http://www.osti.gov/energycitations/product.
biblio.jsp?osti_id=4427744.

Metropolis:1951:TAM

[MR51a] N. Metropolis and George W. Reitwiesner. Table of atomic masses. Technical Report NP-1980, Technical Information Service, Oak Ridge, TN, USA, 1951. 479 pp.

Metropolis:1951:SFT

[MR51b] N. Metropolis and J. R. Reitz. Solutions of the Fermi–Thomas– Dirac equation. Journal of Chemical Physics, 19:555–573, 1951. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

Metropolis:1974:SAA

[MR74] N. Metropolis and Gian-Carlo Rota. Significance arithmetic—on the algebra of binary strings. In Studies in numerical analysis (papers in honour of Cornelius Lanczos on the occasion of his 80th birthday), pages 241–251. Academic Press, New York, NY, USA, 1974.

Metropolis:1978:CSF

[MR78a] N. Metropolis and Gian-Carlo Rota. Combinatorial structure of the faces of the n-cube. SIAM Journal on Applied Mathematics, 35(4):689–694, 1978. CODEN SMJMAP. ISSN 0036-1399 (print),

1095-712X (electronic). URL http://links.jstor.org/sici?
sici=0036-1399%28197812%2935%3A4%3C689%3ACSOTFO%3E2.0.
CO%3B2-U.

Metropolis:1978:LFC

[MR78b] N. Metropolis and Gian-Carlo Rota. On the lattice of faces of the n-cube. Bulletin of the American Mathematical Society, 84(2):284– 286, March 1, 1978. CODEN BAMOAD. ISSN 0002-9904 (print),

1936-881X (electronic).  URL http://projecteuclid.org/
euclid.bams/1183540528;   http://www.osti.gov/energycitations/
product.biblio.jsp?osti_id=5159496&query_id=0.

Metropolis:1983:WVA

[MR83] N. Metropolis and Gian-Carlo Rota. Witt vectors and the algebra of necklaces. Advances in Mathematics, 50(2):95–125, 1983. CO-DEN ADMTA4. ISSN 0001-8708 (print), 1090-2082 (electronic).

Metropolis:1984:CI

[MR84] N. Metropolis and Gian-Carlo Rota. The cyclotomic identity. In Greene [Gre84], pages 19–27. ISBN 0-8218-5029-6 (paperback). ISSN 0271-4132 (print), 1098-3627 (electronic). LCCN QA150 .C647 1984.

Metropolis:1988:SFB

[MR88] N. Metropolis and Gian-Carlo Rota. Symmetric functions: a bijective identity. Proceedings of the American Mathematical Society, 102(1):218–220, 1988. CODEN PAMYAR. ISSN 0002-9939 (print),

1088-6826 (electronic). URL http://links.jstor.org/sici?
sici=0002-9939%28198801%29102%3A1%3C218%3ASFABI%3E2.0.
CO%3B2-X.

Metropolis:1991:SCF

[MR91] N. Metropolis and Gian-Carlo Rota. Symmetry classes: functions of three variables. American Mathematical Monthly, 98(4):328– 332, April 1991. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://links.jstor.org/sici?
sici=0002-9890%28199104%2998%3A4%3C328%3ASCFOTV%3E2.0.
CO%3B2-O.

Metropolis:1993:NEC

[MR93] N. (Nicholas) Metropolis and Gian-Carlo Rota, editors. A New Era in Computation. MIT Press, Cambridge, MA, USA, 1993. ISBN 0-262-63154-7. xii + 241 pp. LCCN QA76.58 .N48 1993. US$13.95.

Metropolis:1952:ESC

[MRR^{+}52] Nicholas Constantine Metropolis, Arianna W. Rosenbluth, Marshall N. Rosenbluth, Augusta H. Teller, and Edward Teller. Equation of state calculations by fast computing machines. Report AECU-2435; LADC-1359, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1952. URL http:/

/www.osti.gov/energycitations/product.biblio.jsp?osti_
id=4390578&query_id=0.

Metropolis:1953:ESC

[MRR^{+}53] Nicholas Metropolis, Arianna W. Rosenbluth, Marshall N. Rosenbluth, Augusta H. Teller, and Edward Teller. Equation of state calculations by fast computing machines. Journal of Chemical Physics, 21(6):1087–1092, June 1953. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL http://link.

aip.org/link/doi/10.1063/1.1699114;   http://pubs.acs.
org/cgi-bin/chemport/version=1.0&coi=1:CAS:528:DyaG3sXltlKhsw%
253D%253D&pissn=0095-2338&pyear=2005&&md5=2eba6ee7d50b361b924c3ff8efeda4b1.

This article introduces the Metropolis algorithm, which the journal Computing in Science and Engineering cited in the top 10 algorithms having the “greatest influence on the development and practice of science and engineering in the 20th Century.” See [BBLS00, BS00], and the Hasting–Metropolis generalization in [Has70]. See also [Bar65, Pes73, Hit03]. According to [Har10, page 263], this paper has been cited more than 10,000 times.

Metropolis:1991:TSC

[MRS91] N. Metropolis, Gian-Carlo Rota, and Joel A. Stein. Theory of symmetry classes. Proceedings of the National Academy of Sciences of the United States of America, 88(19):8415–8419, 1991. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (elec-

tronic). URL http://links.jstor.org/sici?sici=0027-8424%
2819911001%2988%3A19%3C8415%3ATOSC%3E2.0.CO%3B2-V.

Metropolis:1995:SCF

[MRS95] N. Metropolis, Gian-Carlo Rota, and Joel A. Stein. Symmetry classes of functions. Journal of Algebra, 171(3):845–866, 1995. CO-DEN JALGA4. ISSN 0021-8693 (print), 1090-266x (electronic).

Metropolis:1978:PCC

[MRSW78] N. Metropolis, Gian-Carlo Rota, Volker Strehl, and Neil White. Partitions into chains of a class of partially ordered sets. Proceedings of the American Mathematical Society, 71(2):193–196, 1978. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (elec-

tronic). URL http://links.jstor.org/sici?sici=0002-9939%
28197809%2971%3A2%3C193%3APICOAC%3E2.0.CO%3B2-Q.

Metropolis:1973:SAC

[MRT73] N. Metropolis, Gian-Carlo Rota, and S. Tanny. Significance arithmetic: the carrying algorithm. Journal of Combinatorial Theory (Series A), 14:386–421, May 1973. CODEN JCBTA7. ISSN 0097-3165 (print), 1096-0899 (electronic).

Metropolis:1950:STV

[MRvN50] N. C. Metropolis, G. Reitwiesner, and J. von Neumann. Statistical treatment of values of first 2,000 decimal digits of e and of π calculated on the ENIAC. Mathematical Tables and Other Aids to Computation, 4(30):109–111, 1950. CODEN MTTCAS. ISSN 0891-6837.

Metropolis:1967:CMV

[MS67] N. Metropolis and P. R. Stein. On a class of (0, 1) matrices with vanishing determinants. Journal of Combinatorial Theory, 3:191– 198, 1967. CODEN JCTHAR. ISSN 0021-9800 (print), 1878-1756 (electronic).

Metropolis:1970:ESP

[MS70] N. Metropolis and P. R. Stein. An elementary solution to a problem in restricted partitions. Journal of Combinatorial Theory, 9:365– 376, 1970. CODEN JCTHAR. ISSN 0021-9800 (print), 1878-1756 (electronic).

Metropolis:1980:EGP

[MS80] N. Metropolis and P. R. Stein. The enumeration of graphical partitions. European Journal of Combinatorics, 1(2):139–153, 1980. CODEN EJOCDI. ISSN 0195-6698 (print), 1095-9971 (electronic).

Metropolis:1967:SSN

[MSS67] N. Metropolis, M. L. Stein, and P. R. Stein. Stable states of a nonlinear transformation. Numerische Mathematik, 10(1):1–19, May 1967. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Metropolis:1969:PCM

[MSS69] N. Metropolis, M. L. Stein, and P. R. Stein. Permanents of cyclic (0, 1) matrices. Journal of Combinatorial Theory, 7:291– 321, 1969. CODEN JCTHAR. ISSN 0021-9800 (print), 1878-1756 (electronic).

Metropolis:1973:FLS

[MSS73] N. Metropolis, M. L. Stein, and P. R. Stein. On finite limit sets for transformations on the unit interval. Journal of Combinatorial Theory (Series A), 15:25–44, 1973. CODEN JCBTA7. ISSN 0097-3165 (print), 1096-0899 (electronic). See review [May76].

Metropolis:1986:FS

[MSWA86] N. Metropolis, D. H. Sharp, W. J. Worlton, and K. R. Ames, editors. Frontiers of Supercomputing, Los Alamos series in basic and applied sciences 7. University of California Press, Berkeley, CA, USA, 1986. ISBN 0-520-05190-4. LCCN QA76.5 .F76 1983. Papers presented at a conference co-sponsored by the Los Alamos National Laboratory and the National Security Agency, held in Los Alamos on Aug. 15-19, 1983.

Metropolis:1977:SAP

[MT77] N. Metropolis and Stephen M. Tanny. Significance arithmetic: the probability of carrying. Computers and Mathematics with Applications, 3(1):77–81, 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Mengersen:1996:RCH

[MT96] K. L. Mengersen and R. L. Tweedie. Rates of convergence of the Hastings and Metropolis algorithms. Annals of Statistics, 24 (1):101–121, February 1996. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1033066201.

Metropolis:1959:CTI

[MTB59] N. Metropolis, A. L. Turkevich, and R. L. Bivins. Coordinate transformations in intranuclear cascade studies. Technical Report LAMS-2360, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, September 1, 1959. 19 pp. URL http:/

/www.osti.gov/energycitations/product.biblio.jsp?osti_
id=4187884&query_id=0.

Metropolis:1963:PFS

[MTTT63a] N. Metropolis, A. H. Taub, John Todd, and C. B. Tompkins, editors. Experimental arithmetic, high speed computing and mathematics: Proceedings of the fifteenth Symposium in Applied Mathematics of the American Mathematical Society held in Chicago, Illinois, April 12–14, 1962 and Atlantic City, New Jersey, April 16-19, 1962, volume 15. American Mathematical Society, Providence, RI, USA, 1963. LCCN QA297 .S987 1962.

Metropolis:1963:IBM

[MTTT63b] N. C. Metropolis, A. H. Taub, John Todd, and C. B. Tompkins, editors. Interactions between Mathematical Research and High-Speed Computing: Symposia: Selected Papers. American Mathematical Society, Providence, RI, USA, 1963. LCCN ????

Metropolis:1949:MSM

[MU49a] N. C. Metropolis and Stanisław M. Ulam. On motions of systems of mass points randomly distributed on the infinite line. Bulletin of the American Mathematical Society, 55(??):670–671, ???? 1949. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Metropolis:1949:MCM

[MU49b] Nicholas Metropolis and S. Ulam. The Monte Carlo method. Journal of the American Statistical Association, 44(247):335–341, September 1949. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://links.jstor.org/sici?
sici=0162-1459%28194909%2944%3A247%3C335%3ATMCM%3E2.0.
CO%3B2-3;   http://www.jstor.org/stable/2280232.

Metropolis:1952:PRA

[MU52] N. Metropolis and S. Ulam. A property of randomness of an arithmetical function. Report AECU-2038; LADC-1177, Los Alamos Scientific Laboratory, Los Alamos, NM, USA,

1952. 3 pp. URL http://www.osti.gov/energycitations/
product.biblio.jsp?osti_id=4404316&query_id=0. Published

in [MU53].

Metropolis:1953:PRA

[MU53] N. Metropolis and S. Ulam. A property of randomness of an arithmetical function. American Mathematical Monthly, 60(4):252– 253, April 1953. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://links.jstor.org/sici?
sici=0002-9890%28195304%2960%3A4%3C252%3AAPOROA%3E2.0.
CO%3B2-T.

Metropolis:1972:TEH

[MW72] N. Metropolis and J. Worlton. A trilogy on errors in the history of computing. In AFIPS [AFI72], pages 683–691. LCCN QA76 .U2 1972.

Metropolis:1980:TEH

[MW80] N. Metropolis and J. Worlton. A trilogy on errors in the history of computing. Annals of the History of Computing, 2(1):

49–59, January/March 1980. CODEN AHCOE5. ISSN 0164-

1239. URL http://dlib.computer.org/an/books/an1980/pdf/
a1049.pdf;   http://www.computer.org/annals/an1980/a1049abs.

htm. An earlier version is published in [MW72].

Nachbin:1981:MAA

[Nac81] Leopoldo Nachbin, editor. Mathematical analysis and applications: essays dedicated to Laurent Schwartz on the occasion of his 65th birthday, volume 7 of Advances in mathematics. Supplementary studies. Academic Press, New York, NY, USA, 1981. ISBN 0-12-512801-0 (vol. 1), 0-12-512802-9 (vol. 2). LCCN A300 .M294.

Nash:1990:HSC

[Nas90] Stephen G. Nash, editor. A History of Scientific Computing. ACM Press history series. Addison-Wesley and ACM Press, Addison-Wesley and New York, NY 10036, USA, 1990. ISBN 0-201-50814-1. xix + 359 pp. LCCN QA76.17 .H59 1990.

Oppenheimer:1984:US

[OMRS84] J. Robert Oppenheimer, N. (Nicholas) Metropolis, Gian-Carlo Rota, and D. H. (David Howland) Sharp. Uncommon sense. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1984. ISBN 0-8176-3165-8. ix + 195 pp. LCCN Q127.U6 O66 1984.

Peskun:1973:OMC

[Pes73] P. H. Peskun. Optimum Monte–Carlo sampling using Markov chains. Biometrika, 60(3):607–612, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335011.

Ravo:1999:NMM

[Rav99] N. Ravo. Obituary: Nicholas Metropolis: a maker of the A-bomb and computers. New York Times, October 23, 1999. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Metropolis:1959:TJJ

[RBMW59] Manuel Rotenberg, R. Bivins, Nicholas Metropolis, and John W. Wooten, Jr. The 3-j and 6-j symbols. Technology Press, Cambridge, MA, USA, 1959. viii + 498 pp. LCCN QC174.5 .R65. Also published by Crosby Lockwood, London, UK (1959).

Robert:2004:MCS

[RC04] Christian P. Robert and George Casella. Monte Carlo statistical methods. Springer texts in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2004. ISBN 0-387-21239-6. xxx + 645 pp. LCCN QA276

.R575 2004. URL http://www.loc.gov/catdir/enhancements/
fy0818/2004049157-d.html;   http://www.loc.gov/catdir/
enhancements/fy0818/2004049157-t.html.

Richtmyer:1962:CFE

[RDM62] R. D. Richtmyer, Marjorie Devaney, and N. Metropolis. Continued fraction expansions of algebraic numbers. Numerische Mathematik, 4:68–84, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Reitwiesner:1950:EDM

[Rei50] George W. Reitwiesner. An ENIAC determination of π and e to more than 2000 decimal places. Mathematical Tables and Other Aids to Computation, 4(29):11–15, January 1950. CODEN MTTCAS. ISSN 0891-6837.

Richtmyer:1949:MC

[RM49] R. D. Richtmyer and N. C. Metropolis. Modern computing. Physics Today, 2(10):8–15, October 1949. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Roberts:1999:NAR

[Rob99] G. O. Roberts. A note on acceptance rate criteria for CLTs for Metropolis–Hastings algorithms. Journal of Applied Probability, 36(4):1210–1217, December 1999. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3215589.

Rosenbluth:2003:GMC

[Ros03] Marshall Rosenbluth. Genesis of the Monte Carlo algorithm for statistical mechanics. Talk given at Los Alamos National Laboratory., June 9, 2003.

Rota:1986:SCV

[Rot86] Gian-Carlo Rota, editor. Science and Computers: A Volume Dedicated to Nicholas Metropolis, volume 10 of Advances in mathematics. Supplementary studies. Academic Press, New York, NY, USA, 1986. ISBN 0-12-598545-2. xi + 361 pp. LCCN QA76.95 .S36 1986.

Rosenbluth:1954:FRM

[RR54] Marshall N. Rosenbluth and Arianna W. Rosenbluth. Further results on Monte Carlo equations of state. Journal of Chemical Physics, 22(5):881–884, May 1954. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL http://scienze-como.

uninsubria.it/bressanini/montecarlo-history/rosenbluth-
1954.pdf.

Roberts:1996:GCC

[RT96] G. O. Roberts and R. L. Tweedie. Geometric convergence and central limit theorems for multidimensional Hastings and Metropolis algorithms. Biometrika, 83(1):95–110, March 1996. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2337435.

Rudolf:2018:CHR

[RU18] Daniel Rudolf and Mario Ullrich. Comparison of hit-and-run, slice sampler and random walk Metropolis. Journal of Applied Probability, 55(4):1186–1202, December 2018. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic).

URL https://www.cambridge.org/core/journals/journal-
of-applied-probability/article/comparison-of-hitandrun-
slice-sampler-and-random-walk-metropolis/9E32DA02A6E30F0D405C4D1115759BD3.

Richtmyer:1947:SMN

[RUvN47] Robert D. Richtmyer, Stanisław Ulam, and John von Neumann. Statistical methods in neutron diffusion. Technical Report LAMS-551, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, April 9, 1947. 22 pp. Republished in typeset form in [Hur85].

S:1969:BRQ

[S.69] D. S. Book review: Square Roots of Integers 2 to 15 in Various Bases 2 to 10: 88062 Binary Digits or Equivalent by W. A. Beyer, N. Metropolis, and J. R. Neergaard. Mathematics of Computation, 23(107):679, July 1969. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://links.jstor.org/sici?sici=0025-5718%
28196907%2923%3A107%3C679%3ASROI2T%3E2.0.CO%3B2-C.

Sasaki:1991:EDS

[Sas91] Galen Sasaki. The effect of the density of states on the Metropolis algorithm. Information Processing Letters, 37(3):159–163, February 18, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Segre:1976:PSN

[Seg76] Emilio Segrè. Personaggi e scoperte nella fisica contemporanea. (Italian) [Personalities and discoveries in contemporary physics]. Biblioteca della EST. Edizioni scientifiche e tecniche Mondadori, Milano, Italia, 1976. 297 pp. LCCN QC7 .S44. Ciclo di lezioni tenute dal nov. 1972 fino al marzo 1973.

Segre:1980:XRQ

[Seg80] Emilio Segrè. From X-rays to quarks: modern physicists and their discoveries. W. H. Freeman, New York, NY, USA, 1980. ISBN 0-7167-1146-X, 0-7167-1147-8 (paperback). ix + 337 pp. LCCN QC7 .S4413. English translation of [Seg76].

Segre:1983:PSN

[Seg83] Emilio Segrè. Personaggi e scoperte nella fisica contemporanea: dai raggi X ai quark. Biblioteca della EST, 0303-2752. Edizioni scientifiche e tecniche Mondadori (IS), Milano, Italia, second edition, 1983. 297 pp. LCCN ????

Segre:1984:PML

[Seg84] Emilio Segrè. Les physiciens modernes et leurs découvertes: des rayons X aux quarks. (French) [Modern physicists and their discoveries: from X-rays to quarks]. Temps des sciences. Fayard, Paris, France, 1984. 456 pp. LCCN QC7 .S4414 1984. French translation of [Seg76].

Segre:1986:MKR

[Seg86a] Emilio Segre. Mi-karne Ranotgen ove-ad kvarkim: fisikaim modernim ove-tagliyotehe. Keter, Yerushalayim, Israel, 1986. 331 pp. LCCN ???? Hebrew translation of [Seg76].

Segre:1986:WLM

[Seg86b] Emilio Segrè. Wu li ming ren he wu li fa xian. Zhi shi chu ban she, Shanghai, People’s Republic of China, 1986. iii + 366 pp. LCCN QC7 .S4412. Mandarin Chinese translation by Zuwei Liu of [Seg76].

Segre:1987:RXA

[Seg87] Emilio Segre. Dos raios X aos quarks: fisicos modernos e suas descobertas. (Portuguese) [From X-rays to quarks: modern physicists and their discoveries], volume 24 of Pensamento cientifico. Ed. UnB, Brasilia, Brazil, 1987. ISBN 85-230-0078-X. 345 pp. LCCN ????

Segre:2007:XRQ

[Seg07] Emilio Segrè. From X-rays to quarks: modern physicists and their discoveries. Dover classics of science and mathematics. Dover Publications, Inc., New York, NY, USA, 2007. ISBN 0-486-45783-4. ix + 339 pp. LCCN QC7 .S4413 2007. URL http://www.loc.gov/

catdir/enhancements/fy0702/2006102450-d.html.

Shanks:1971:CSS

[Sha71] Daniel Shanks. Corrigendum to: “Statistical study of digits of some square roots of integers in various bases” (Math. Comp. 24 (1970), 455–473) by W. A. Beyer, N. Metropolis and J. R. Neergaard. Mathematics of Computation, 25(114):409, 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.org/journals/mcom/1971-25-
114/S0025-5718-1971-0400639-3.

Shortley:1960:BRJ

[Sho60] George Shortley. Book review: The 3-j and 6-j Symbols, by Manuel Rotenberg, R. Bivins, N. Metropolis, and John K. Wooten, Jr. Mathematics of Computation, 14(72):382–383, 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842

(electronic). URL http://links.jstor.org/sici?sici=0025-
5718%28196010%2914%3A72%3C382%3ATAS%3E2.0.CO%3B2-C. See

[RBMW59].

Smillie:1982:RMA

[SR82] K. W. Smillie and Saul Rosen. Reviews: N. Metropolis, et al.: A History of Computing in the Twentieth Century; R. L. Wexelblat: History of Programming Languages; T. Forester: The Microelectronics Revolution; capsule reviews. Annals of the History of Computing, 4(1):69–78, January/March 1982. CODEN AH-

COE5. ISSN 0164-1239. URL http://dlib.computer.org/an/
books/an1982/pdf/a1069.pdf;   http://www.computer.org/
annals/an1982/a1069abs.htm.

Sofroniou:2005:PNC

[SS05] Mark Sofroniou and Giulia Spaletta. Precise numerical computation. Journal of Logic and Algebraic Programming, 64 (1):113–134, July 2005. ISSN 1567-8326 (print), 1873-5940

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1567832604000785.

Stern:1981:BRH

[Ste81] Nancy Stern. Book review: A History of Computing in the Twentieth Century, by N. Metropolis, J. Howlett, and Gian-Carlo Rota. Science (New Series), 212(4494):536–537, May 1981. ISSN 0036-8075 (print), 1095-9203 (electronic). URL

http://links.jstor.org/sici?sici=0036-8075%2819810501%
293%3A212%3A4494%3C536%3ACWR%3E2.0.CO%3B2-6.

Stapp:1957:PSA

[SYM57] H. P. Stapp, T. J. Ypsilantis, and N. Metropolis. Phaseshift analysis of 310-Mev proton-proton scattering experiments. Physical Review, 105(1):302–310, January 1, 1957. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://prola.aps.org/pdf/PR/v105/i1/p302_1;
 http://www.osti.gov/energycitations/product.biblio.jsp?
osti_id=4381983&query_id=0.

Taub:1961:JNCa

[Tau61a] A. H. Taub, editor. John von Neumann: Collected Works: Volume I: Logic, Theory of Sets and Quantum Mechanics. Pergamon Press, New York, NY, USA, 1961. x + 654 pp. LCCN ???? See also volumes II–VI [Tau61b, Tau63a, Tau62, Tau63b, Tau63c].

Taub:1961:JNCb

[Tau61b] A. H. Taub, editor. John von Neumann: Collected Works. Volume II: Operators, Ergodic Theory and Almost Periodic Functions in a Group. Pergamon Press, New York, NY, USA, 1961. x + 568 pp. LCCN ???? See also volumes I, III–VI [Tau61a, Tau63a, Tau62, Tau63b, Tau63c].

Taub:1962:JNC

[Tau62] A. H. Taub, editor. John von Neumann: Collected Works. Volume IV: Continuous Geometry and Other Topics. Pergamon Press, New York, NY, USA, 1962. x + 516 pp. LCCN ???? See also volumes I–III, V–VI [Tau61a, Tau61b, Tau63a, Tau63b, Tau63c].

Taub:1961:JNCc

[Tau63a] A. H. Taub, editor. John von Neumann: Collected Works. Volume III: Rings of Operators. Pergamon Press, New York, NY, USA, 1961–1963. ix + 574 pp. LCCN ???? See also volumes I–II, IV–VI [Tau61a, Tau61b, Tau62, Tau63b, Tau63c].

Taub:1963:JNCa

[Tau63b] A. H. Taub, editor. John von Neumann: Collected Works. Volume V: Design of Computers, Theory of Automata and Numerical Analysis. Pergamon Press, New York, NY, USA, 1963. ix + 784 pp. LCCN ???? See also volumes I–IV, VI [Tau61a, Tau61b, Tau63a, Tau62, Tau63c].

Taub:1963:JNCb

[Tau63c] A. H. Taub, editor. John von Neumann: Collected Works. Volume VI: Theory of Games, Astrophysics, Hydrodynamics and Meteorology. Pergamon Press, New York, NY, USA, 1963. x + 538 pp. LCCN ???? See also volumes I–V [Tau61a, Tau61b, Tau63a, Tau62, Tau63b].

Teller:1955:WMP

[Tel55] Edward Teller. The work of many people. Science, 121(3139): 267–275, February 25, 1955. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/

stable/1681529.

Tsuda:1995:NMM

[TFY95] N. Tsuda, A. Fujitsu, and T. Yukawa. Note on the Metropolis Monte Carlo method on random lattices. Computer Physics Communications, 87(3):372–374, June 1995. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/001046559400129P.

Tierney:1994:MCE

[Tie94] Luke Tierney. Markov chains for exploring posterior distributions. Annals of Statistics, 22(4):1701–1728, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176325750;   http://www.jstor.org/stable/2242477.

Tomayko:1989:ACA

[TMRL89] James E. Tomayko, Nicholas C. Metropolis, R. D. Richtmyer, and John A. N. Lee. Anecdotes: The case against automatic programming. Annals of the History of Computing, 11(4):322–326, October/December 1989. CODEN AH-

COE5. ISSN 0164-1239. URL http://dlib.computer.org/an/
books/an1989/pdf/a4322.pdf;   http://www.computer.org/
annals/an1989/a4322abs.htm.

Tropp:1983:BRH

[Tro83] Henry S. Tropp. Book review: A history of computing in the twentieth century: Edited by N. Metropolis, J. Howlett, and Gian-Carlo Rota. New York (Academic Press). 1980. xix + 659 pp. $29.50. Historia Mathematica, 10(2):233–236, May 1983. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0315086083900721.

VanDerwerken:2017:EGS

[Van17] Douglas VanDerwerken. Not every Gibbs sampler is a special case of the Metropolis–Hastings algorithm. Communications in Statistics: Theory and Methods, 46(20):10005–10009, 2017. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/03610926.2016.1228961.

Waterman:2004:NG

[Wat04a] Michael Waterman. Nick the Greek. In Skiing the Sun: New Mexico Essays [Wat04b], pages 17–20. URL http://www-hto.

usc.edu/people/msw/newmex.pdf. Unpublished manuscript.

Waterman:2004:SSN

[Wat04b] Michael Waterman. Skiing the sun: New Mexico essays. Web doc-

ument, July 21, 2004. URL http://www-hto.usc.edu/people/

msw/newmex.pdf. Unpublished manuscript.

Weiss:1988:BOP

[Wei88] Eric A. Weiss. Biographies: Oh, pioneers! Annals of the History of Computing, 10(4):348–361, October/December 1988. CODEN AH-

COE5. ISSN 0164-1239. URL http://dlib.computer.org/an/
books/an1988/pdf/a4348.pdf;   http://www.computer.org/
annals/an1988/a4348abs.htm.

Wu:2005:EMA

[WW05] Samuel S. Wu and Martin T. Wells. An extension of the Metropolis algorithm. Communications in Statistics: Theory and Methods, 34 (3):585–596, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yost:2002:BGR

[Yos02] Jeffrey R. Yost. A Bibliographic Guide to Resources in Scientific Computing, 1945–1975, volume 15 of Bibliographies and indexes in library and information science, 0742-6879. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 2002. ISBN 0-313-31681-3. 263 pp. LCCN Z7405.D37 Y67 2002; Q183.9. URL

http://www.loc.gov/catdir/toc/fy035/2002069622.html.